### v1

# Conceptualization of this study, Methodology, Writing

Abstract
Understanding actions within surgical workflows is essential for evaluating post-operative outcomes. However, capturing long sequences of actions performed in surgical settings poses challenges, as individual surgeons have their unique approaches shaped by their expertise, leading to significant variability. To tackle this complex problem, we focused on segmentation with precise boundaries, a demanding task due to the inherent variability in action durations and the subtle transitions often observed in untrimmed videos. These transitions, marked by ambiguous starting and ending points, complicate the segmentation process. Traditional models, such as MS-TCN, which depend on large receptive fields, frequently face challenges of over-segmentation (resulting in fragmented segments) or under-segmentation (merging distinct actions). Both of these issues negatively impact the quality of segmentation. To overcome these challenges, we present the Multi-Stage Boundary-Aware Transformer Network (MSBATN) with hierarchical sliding window attention, designed to enhance action segmentation. Our proposed approach incorporates a novel unified loss function that treats action classification and boundary detection as distinct yet interdependent tasks. Unlike traditional binary boundary detection methods, our boundary voting mechanism accurately identifies start and end points by leveraging contextual information. Extensive experiments using three challenging surgical datasets demonstrate the superior performance of the proposed method, achieving state-of-the-art results in F1 scores at thresholds of 25% and 50%, while also delivering comparable performance in other metrics.
keywords:
[…]
1 Introduction
Surgical action segmentation involves analyzing complex elements such as instruments, organs, fluids, or smoke to identify and categorize distinct actions or phases in surgical videos Mascagni et al. (2022). This task is crucial for enhancing surgical training by providing detailed feedback to trainees, improving workflow efficiency and patient safety through real-time monitoring and optimization of procedures, and advancing the performance of autonomous or semi-autonomous surgical robots capable of understanding and executing complex tasks. Recent advancements in this domain have led to the creation of sophisticated architectures tailored to the unique challenges of surgical environments (like Gazebo, ROS2 software development kit (SDK), etc), enabling better decision-making services during operations and driving progress in surgical robotics and education Kiyasseh et al. (2023).
Surgical videos are often heavily annotated manually, where each frame is labelled to a surgical phase or sub-phase or specific action/gesture, as well as the surgical instruments present in that frame Psychogyios et al. (2023); Twinanda et al. (2016). Recent datasets have incorporated action triplets Nwoye et al. (2022), which include action, instrument, and organ information for frame-wise action recognition and instruments or organs segmentation Nwoye et al. (2023); Chadebecq et al. (2023). However, temporal evolution hierarchies are not well defined as they are subjective in nature. The high to low-level temporal component of a surgical operation or intraoperative activity is termed phase, steps and tasks. For more details please refer attached supplementary material file.
Figure 1 demonstrates an effective approach to handle temporal variations by modelling start-end transition in long videos. It starts with frame-wise action and boundary recognition, and then it goes through a segment-wise robust refinement process to avoid over-under segmentation.
For efficient analysis of surgical workflows and improving postoperative outcomes, a robust understanding of both spatial context and temporal relationships, the arrangement and transitions between actions or gestures, are essential Chadebecq et al. (2023). In this regard, most publicly available datasets often lack these temporal hierarchies and their evolutions. Additionally, similarities between different phases, including unclear or ambiguous starting, ending and transitions between actions 2, lead to inaccurate recognition.
Figure 1: Overview of the proposed framework, which focuses on segments encompassing the action’s start, centre, and end, enabling the model to capture nuanced transitions and improve boundary precision.
For improved surgical video analysis, most existing models are built based on deep neural networks which are broadly divided into single-stage and multi-stage where the architectures evolved from RNN to Temporal Convolution Network(TCN)Farha and Gall (2019); Li et al. (2020); Czempiel et al. (2020); Yi et al. (2022), and Transformer Yang et al. (2024); Zhang et al. (2022); Ding and Li (2022); Nwoye et al. (2022); Shi et al. (2023) based approaches. Single-stage models make predictions directly from visual-temporal features. These models utilize a small window of time around the target frame to capture spatial-temporal features, which minimizes computational costs. Early models like EndoNetTwinanda et al. (2016); Twinanda (2017) focused on detecting tool presence and recognizing phases using RNN-based methods. EndoNetTwinanda et al. (2016) is the first method proposed for two tasks simultaneously in a multitasking manner, including tool presence detection and phase recognition. EndoLSTM Twinanda (2017), as the improved EndoNet, replaced hierarchical hidden Markov models (HHMMs) with LSTM for temporal refinement and improved the results substantially. Besides, in Mondal et al. (2019), a multitask deep learning framework was proposed comprising ResNet-50 and Bi-LSTM. It performs better in tool detection and phase recognition. In Jin et al. (2020), a single-staged multitask recurrent convolutional network was proposed with correlation loss (MTRCNet-CL) for surgical video analysis. This multitask learning method used a new correlation loss to minimize tool and phase prediction divergences.
However, RNN and LSTM-based approaches often struggle to capture long-term dependencies that are essential for accurately modelling the entire surgical process. For instance, models like Actionformer Zhang et al. (2022) and Tridet Shi et al. (2023) integrate Transformer architecture for action recognition. However, these single-stage models are designed for small actions while each phase or action segment in surgery takes several minutes, or the whole procedure sometimes takes more than 1 hour. Because of this research, the signal-stage models are not ideal for long untrimmed video analysis.
Multi-stage models, on the other hand, include an additional refinement stage to improve the initial predictions by correcting the errors. Multi-stage models are considered more suitable for surgical phase recognition to optimize the complexity and ordered nature of surgical activitiesCzempiel et al. (2020); Van Amsterdam et al. (2022). Multi-stage models such as TeCNOCzempiel et al. (2020) and Farha and Gall (2019); Li et al. (2020) employ multiple temporal convolutional networks to refine predictions in subsequent stages. However, these methods often neglect high-level semantic information and struggle to refine predictions effectively in an end-to-end manner, leading to over or under-segmentation errors. This is because the refinement stage focuses on adjusting initial predictions without fully understanding the actions involved during training, and errors remain during inference. This can lead to misinterpretations in complex scenarios where context is crucial. Recently, in Ramesh et al. (2021), a multitask multistage temporal convolutional network (MTMS-TCN) was proposed to jointly learn the tasks of phase and step recognition. Some studies applied sensor data with attention mechanism to enhance the understanding of surgical processes Van Amsterdam et al. (2022); Goldbraikh et al. (2023), however aligning sensor(kinematic) data with corresponding video frames is complex due to differences in sampling rates and potential time lags(different frequencies), leading to misalignment issues. To improve the segmentation, some research has focused on incorporating segment-level modeling with transformer Ding and Li (2022); Jin et al. (2022). Their main idea is to use high-level segments from surgical videos to clarify ambiguous semantics presented in neighboring frames.
Figure 2: Ambiguous action ordering in untrimmed long surgical videos
To overcome the above limitations, the current methods use sliding window detectors Czempiel et al. (2020), but they struggle to capture long-term patterns effectively Twinanda et al. (2016); Twinanda (2017); Mondal et al. (2019); Czempiel et al. (2020). Additionally, these approaches overlook the finer details that are essential for understanding the intricate relationships between spatial and temporal features Kiyasseh et al. (2023), resulting in over and under-segmentation errors. The over-segmentation occurs when a segment is incorrectly divided into multiple action segments, leading to the detection of more actions than actually exist. These approaches are overly sensitive to minor variations in the input data, causing them to interpret slight changes as new action boundaries Ishikawa et al. (2021). On the other hand, Under-segmentation happens when distinct actions are mistakenly merged into a single segment. This typically arises due to a lack of sufficient temporal resolution or failure to recognize subtle transitions between different actions. So, the model overlooks actual action boundaries, resulting in fewer detected action segments than actually exist.
Most of the approaches fail to model the continuous and smooth transitions characteristic of surgical activities due to treating actions and boundaries as discrete events without accounting for their gradual shifts between actions Fard et al. (2017); Ishikawa et al. (2021). We noticed that the self-attention mechanism has difficulty focusing on important local areas, particularly at the initial layers. Additionally, the encoder-decoder architectures do not effectively refine predictions based on the temporal relationships between actions Yi et al. (2021). To address these issues, first, we propose a framework that extracts spatial features with DinoV2 by discarding irrelevant large patch tokens with register tokens Oquab et al. (2023). Second, for temporal modelling, we present an encoder-decoder model that is built based on the combination of dilated TCN, unified hierarchical and sliding window attention, which captures coarse-to-fine context. The hierarchical representation with dynamic sliding window attention is used to handle long input sequences that organize the attention layers hierarchically to first focus on local relationships and then gradually capture global dependencies. Last, we develop a unified loss function by combining the loss functions of individual tasks into a single action segmentation to achieve effective boundary detection. This holistic approach allows one to learn multiple tasks simultaneously, leveraging shared representations to improve overall segmentation performance.
Our approach introduces the following key innovations:
  • We present a multi-stage TCN-Transformer architecture with dilated acausal convolution to widen the receptive field that predicts and refines both actions and boundaries.
  • We introduce hierarchical sliding window attention mechanisms (HSWA) to resolve ambiguities among action orders from different temporal resolution. It links high-level action segment structures with fine-grained frame-wise details, enabling the model to better understand long-term temporal dependencies and immediate frame-level changes.
  • We propose a boundary aware loss function that gradually measures cosine similarity from center to boundaries inside action segments that assist in minimizing over or under-segmentation error.
The remainder of this article is structured as follows: Section 2 focuses on related work. Section 3 describes the proposed framework. The model performance is discussed based on experimental evaluation results in Section 4, and the article concludes with Section 5.
2 Related Work
2.1 Action Recognition and Segmentation
For the surgical domain, EndoNetTwinanda et al. (2016) was developed for phase recognition in cholecystectomy videos by learning inherent visual features through a CNN architecture. Recognizing the importance of tool information in phase recognition, EndoNet adopted a multi-task framework that incorporates tool presence detection into the feature learning process. An extension of AlexNet was designed with dual objectives: one part focused on detecting the presence of surgical tools, while the other was responsible for recognizing seven distinct surgical phases. To capture temporal dependencies, EndoNet integrated a Hierarchical Hidden Markov Model (HMM), enabling the modelling of temporal transitions between phases. The architecture is further improved by LSTMTwinanda (2017) and ResnetTwinanda et al. (2018). However, these methods struggle with accurately identifying transition frames between action phases. To overcome this limitation, a sequence-to-sequence model with an attention mechanism was proposed Namazi et al. (2019); Yi et al. (2021), which accommodates variable-length inputs and outputs, making it well-suited for detecting overlapping or sequential surgical tasks.
Temporal Action Detection (TAD) methodologies are mainly classified into two-stage and one-stage-based methods. The two-stage approach, inspired by object detection frameworks like Faster R-CNN, involves an initial phase of generating action proposals, followed by subsequent stages of classification and refinement Chao et al. (2018); Keisham et al. (2022). Although this method often achieves higher accuracy due to its detailed localization process, it tends to be computationally intensive, posing challenges for real-time applications Yi et al. (2022). On the other hand, the one-stage approach together consolidates action proposal generation and classification in a single framework, which offers computational efficiency for real-time scenarios. For instance, models like ActionFormer Zhang et al. (2022), and Surgformer Yang et al. (2024) are transformer-based architectures that have achieved enhanced precision of action localization while maintaining computational efficiency. These models capture long-range dependencies and contextual information that helps improve the accuracy of action segmentation and detection. However, one-stage methods have historically faced difficulties in precisely localizing action boundaries in long untrimmed videos due to their small receptive fields, which is crucial for optimal performance in TAD.
2.2 Boundary Detection
Identifying action boundaries within untrimmed videos poses unique challenges due to the inherent ambiguity in determining precise start and end points of actions. Initial action boundary detection approaches used probabilistic models to capture temporal dynamics. For example, a spatiotemporal Gaussian Mixture Model (GMM) with Variational Bayesian (VB) Loukas et al. (2016) inference was employed to dynamically estimate the number of components in the GMM, which eliminates the need to predefine the number of clusters. VB inference approximated the full posterior distribution, which improved model flexibility and reduced redundant components. Label assignment and tracking were achieved through Kullback-Leibler (KL) distance, ensuring consistent tracking across video clips. In contrast to rigid boundary detection methods, the Soft-Boundary Unsupervised Gesture Segmentation (Soft-UGS) Fard et al. (2017) approach was proposed as a soft-boundary clustering technique that allowed for gradual transitions between actions. This method utilized kinematic trajectory data, Dynamic Time Warping (DTW), and probabilistic clustering to segment surgical gestures effectively in real-time applications. While these probabilistic models laid the groundwork for more advanced methods, they had limitations in precision, particularly in detecting abrupt transitions in video content.
On the other hand, the advent of deep learning introduced new techniques for addressing the temporal and spatial complexities of action boundary detection. For the surgical domain, sequence models were applied to map variable-length input sequences to output sequences while incorporating attention mechanisms for more accurate boundary localization. A double attention mechanism(such as Namazi et al. (2019)) was introduced to focus specifically on the beginning and end of each surgical phase, offering a refined approach to boundary detection in complex tasks like surgery.
The above methods for temporal segmentation face the issue of over-segmentation because they segment actions into too many sub-components, leading to inaccurate predictions. The Action Segment Refinement Framework (ASRF)Ishikawa et al. (2021) was introduced to address this problem by decoupling frame-wise classification from boundary regression. ASRF refines action segmentation through two branches: one for action predictions and another for boundary regression. By focusing specifically on predicted action boundaries, ASRF achieved overall action localization accuracy. ASRF*Yi et al. (2021) improved the performance with an attention mechanism.
Recent methods, such as TriDetShi et al. (2023) and Elastic Moment Bounding (EMB)Huang et al. (2022), have further extended the capabilities of one-stage detection for real-time applications. TriDet introduced the Trident-head detection mechanism, which models action boundaries by estimating relative probability distributions instead of directly predicting offsets. This innovation significantly improved boundary localization accuracy while reducing computational overhead. EMB tackles uncertainties in temporal boundaries by transforming rigid and manually labelled endpoints into an elastic set of boundaries; however, it does tend to generate a high number of false positives.
Building upon these advancements, our action boundary network employs a voting mechanism to improve segmentation by accurately finding the start and end of an action. By utilizing the predicted action boundaries, our framework refines frame-wise predictions, ultimately improving the performance of action segmentation.
Table 1: Notation table
Symbol Description
$T$ Sequence length (number of frames).
$D$ Dimensionality of frame features.
$\mathbf{x}_{t}$ Feature vector of the $t$-th frame.
$C$ Set of action categories (including background).
$S$ Set of action segments.
$W$ Base window size for sliding attention.
$d$ Dilation factor for dilated attention.
$T_{s}$ Sequence length at scale $s$.
$\alpha_{ij}^{s}$
Attention weight for token $x_{i}$ attending
to $x_{j}$ at scale $T_{s}$.
$\mathcal{L}_{\text{dice}}$ Dice loss for segment overlap.
$\mathcal{L}_{\text{sim}}$ Gaussian-weighted cosine similarity loss.
$\mathcal{L}_{\text{boundary}}$ Boundary-aware loss for segment precision.
$\alpha,\beta,\gamma,\delta$ Weights for different loss components.
$G(t,t+1)$ Gaussian weight emphasizing temporal proximity.
$G_{\text{boundary}}(t)$ Gaussian weight centered on segment boundaries.
3 Proposed Framework
The action segmentation task involves predicting the action category for each frame in a video. Formally, given a sequence of $T$-frame features $X=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times d},\quad\text{where }\mathbf{x}_{t}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature of the $t$-th frame. The objective is to predict a sequence of actions $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{T}],\quad\text{where }\hat{y}_{t}\in C.$ The set $C=\{0,1,2,\ldots,C,-100\}$ includes all action categories in the dataset, with $-100$ representing the background class. The frame-wise ground truth labels are $Y=[y_{1},y_{2},\ldots,y_{T}],$ accompanied by segment information $S=\{(y_{s1},y_{e1},a_{c1}),\ldots,(y_{sn},y_{en},a_{cn})\},$ where $y_{s}$ and $y_{e}$ denote the start and end of a segment, $a_{c}$ represents the action class label, and $n$ is the number of segments, including background segments and notations are listed in Table 1. The goal is to accurately predict action transitions and segment boundaries. Since action transitions are continuous and boundaries are often subjective, considering this 5% of the total segment length is labeled as a transition buffer zone at the start and end of each segment.
Our proposed framework is depicted in figure 3 that enables three distinct parts. First, the frame wise feature extraction, second, Encoder-Decoder architecture with hierarchical sliding window attention, and last, boundary aware loss function that guides to detect the transition with start-center-end points.
Figure 3: Proposed Multi-Stage Boundary-Aware Transformer Network (MSBATN) architecture
3.1 Frame-Wise Feature Extraction
Dino with Register Oquab et al. (2023) (Dinov2-Reg) captures complex spatial features in a self-supervised manner, Figure 4 depicts the performance in comparison with ViT and Resnet.
Figure 4: Comparison of frame feature extractors.
Dinov2-Reg model extracts features using a self-supervised vision transformer model, which processes image-like inputs (frames in this case) to output dense, high-dimensional feature representations. The input feature dimensions (256, 1536), where 256 denotes the 16x16 patch dimension, and 1536 is the feature dimension. A 2D convolutional layer with a kernel size of (256, 1536) is applied to the input frames, which projects the high-dimensional DINO features (256x1536) into a lower-dimensional representation (1x2048). This is equivalent to performing dimensionality reduction via a convolutional operation, which retains important information while reducing the computational complexity for further processing in a Multi-Stage Boundary-Aware Transformer Network (MSBATN). For more details, please refer to the attached supplementary material file.
3.2 Multi-Stage Boundary-Aware Transformer Network (MSBATN)
We present MSBATN transformer-based encoder-decoder architecture without positional encoding to improve the performance of the segmentation task. Each encoder-decoder block comprises a series of TCN and hierarchical sliding window attention modules.
3.2.1 Encoder-Decoder
The frame-wise features $(T,D)$ are passed through an encoder block. The encoder block comprises a series of dilated Acausal Temporal Convolution feed-forward layer blocks to capture context from past and future, followed by the hierarchical sliding window attention module and, lastly, project the encoded features. Instead of using traditional fully connected layers, dilated temporal convolutions are employed to expand the receptive field without increasing the number of parameters or sacrificing resolution. It introduces local inductive biases to effectively capture temporal dependencies and patterns over extended sequences. The model adopts a hierarchical pattern in which the attention layers simultaneously focus on local-global information, mirroring techniques utilized in longformerBeltagy et al. (2020). To capture the global context, we calculated the hierarchical attention at different scales, and then we aggregated to obtain the global context. A detailed discussion is added in the next subsection.
The decoder considered the temporal relationships between action-boundary segments to improve the initial predictions. Similar to the encoder block, the decoder begins with a temporal convolutions block, attention module and output. However, instead of acausal convolution, it has causal convolution, so it does not have any information about the future. Each attention block uses a weighted residual connection to combine the outputs from the feed-forward layers, with the weight $\alpha$ decreasing exponentially across decoders to prevent error accumulation. Last, the decoder generates action boundary labels. The network is comprised of one encoder and three decoders, with each encoder and decoder containing ten TCN-attention blocks. We set the feature dimension in the encoder and decoder layer to 256. Additionally, a temporal dropout layer is applied to the input feature of the encoder, which randomly drops the entire feature channel with a dropout rate of 0.3 to mitigate the over-fitting issue.
3.2.2 Hierarchical dual Sliding Window Attention
Instead of computing attention scores across all token pairs, as in self-attention, we employed a sparse-attention mechanism. This approach uses a hierarchical sliding window that restricts attention to a fixed-size window around each token at a specific layer. The window slides and aggregates scores, meaning each token attends only to its neighboring tokens within the specified window. This reduces both space and computational complexity to $O(n)$. To capture temporal dependencies at varying distances, a dilated attention window is applied with increasing dilation rates $(0,1,2,3,\dots)$ in each layer. This approach allows for capturing both local details from individual frames and long-term temporal patterns across the sequence. The action head and boundary head predict the action and start-end labels, respectively Nwoye et al. (2022). For a sequence length $T$ and we define the base temporal window size $64$, smallest window size $16$, largest window size $512$, average number of action segments $S$. To expand the receptive field, the effective window size $W_{d}$ for each layer $d$ is calculated using dilation as $W_{d}=W\cdot(1+d)$ where $W$ is the base window size, $d$ is the dilation factor ($d=0,1,2,\dots$). For Dual Window Sliding Attention, the effective receptive field starts at $16$ and grows up to $512$, that means $W_{d}\in[16,512]$. Local attention is computed within each window and then aggregated across windows. The temporal hierarchical attention operates at multiple scales. The scale $T_{s}$ is reduced iteratively $T_{s}=\frac{T}{2^{s}}$ for $s=0,1,2,\dots$, until $T_{s}\geq 512$. The average number of scales is approximately $\log_{2}\left(\frac{T}{S}\right)$, based on the sequence length $T$ and the average number of action segments $S$. At each scale $T_{s}$, global attention weights are computed using a sparse mechanism with dilation. For a token $x_{i}$ attending to token $x_{j}$ at scale $T_{s}$, the attention weight $\alpha_{ij}^{s}$ is
$\alpha_{ij}^{s}=\frac{\exp\left(e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}^{s}}\exp\left(e_{ik}^{s}\right)}$ (1)
where $e_{ij}^{s}$ is the attention score at scale $T_{s}$, $\mathcal{N}_{i}^{s}$ is the neighborhood of token $x_{i}$ at scale $T_{s}$, defined by the dilated window size $W_{d}$. The final attention is calculated by aggregating scores across all scales and normalizing them with a softmax function:
$\alpha_{ij}=\frac{\exp\left(\sum_{s}e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\sum_{s}e_{ik}^{s}\right)}$ (2)
where $\mathcal{N}_{i}=\bigcup_{s\in S}\mathcal{N}_{i}^{s}$ is the union of neighborhoods across all scales, $\alpha_{ij}$ is the final attention weight for token $x_{i}$ attending to $x_{j}$. At each scale $T_{s}$, the attention score $e_{ij}^{s}$ between tokens $x_{i}$ and $x_{j}$ within the neighborhood $\mathcal{N}_{i}^{s}$ is:
$e_{ij}^{s}=\text{sim}(x_{i},x_{j}^{s})$ (3)
where $x_{j}^{s}$ is the representation of token $x_{j}$ at scale $s$, $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product). By restricting computations to sparse attention neighborhoods $\mathcal{N}_{i}$, the computational overhead is significantly reduced while retaining efficiency. The number of scales is determined byNumber of $\text{Number of scales}=\log_{2}\left(\frac{T}{S}\right).$. The scale reduction follows $T_{s}=\frac{T}{2^{s}},\quad s=0,1,2,\dots,\log_{2}\left(\frac{T}{S}\right),$ reducing the sequence length $T$ to approximately $\frac{T}{S}$. For example, if $T=1000$ and $S=4$, the segment sizes would be $(1000,500,250,125,\dots)$.
3.3 Boundary-aware Segmentation Loss
Figure 5: Segment-wise Gaussian-cosine similarity scores for SAR-RARP
In temporal action segmentation, accurately capturing action transitions and segment-level dynamics across varying time scales is crucial. Surgical videos are typically long, and the duration of the segment fluctuates significantly, as shown in Table 2. To address this, we propose a Boundary-Aware Segmentation Loss that integrates multiple loss functions, each targeting specific aspects of the segmentation task. To handle class imbalances and ensure precise classification at each timestep, we employ Focal Loss. This loss function down-weights well-classified examples, allowing the model to focus on harder, misclassified examples. This approach is particularly beneficial in datasets where certain action classes are underrepresented. To enhance the overlap between predicted and ground truth segments, we utilize Dice Loss derived from the Dice Coefficient:
Dice $\text{Dice Coefficient}=\frac{2\times|A\cap B|}{|A|+|B|}$ (4)
where $A$ and $B$ represent the predicted and ground truth sets, respectively. The Dice Loss is then formulated as:
$\mathcal{L}_{\text{dice}}=1-\text{Dice Coefficient}$ (5)
This formulation emphasizes the overlap between predictions and ground truth, making it sensitive to both false positives and false negatives. However, focal and dice loss primarily focuses on frame-wise classification accuracy without explicitly considering the temporal relationships between frames. As can be seen in figure 5, the uppermost figure shows the similarity between neighbouring frames, which are ambiguous; we smoothen the distribution by measuring centre-boundary Gaussian weighted cosine similarity depicted in the middle of the figure. To capture temporal dependencies and ensure smooth transitions within action segments, we introduce a loss term based on Gaussian-weighted cosine similarity. This loss penalizes deviations in feature representations between consecutive frames, weighted by a Gaussian function centred on action segments:
$\mathcal{L}_{\text{sim}}=\sum_{t=1}^{T-1}G(t,t+1)\left(1-\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})\right)$ (6)
where $G(t,t+1)$ is the Gaussian weight emphasizing temporal proximity, and $\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})$ denotes the cosine similarity between feature vectors at frames $t$ and $t+1$. To further refine boundary precision, we incorporate a Gaussian Similarity Truncated Mean Squared Error (MSE) loss. This loss focuses on boundary regions, applying a Gaussian weight to the MSE, which emphasizes errors near segment boundaries:
$\mathcal{L}_{\text{boundary}}=\sum_{t=1}^{T}G_{\text{boundary}}(t)\cdot\min\left((\hat{b}_{t}-b_{t})^{2},\tau\right)$ (7)
Where $G_{\text{boundary}}(t)$ is a Gaussian weight centred on the boundary, $\hat{b}_{t}$ and $b_{t}$ are the predicted and ground truth boundary positions at frame $t$, and $\tau$ is a truncation threshold to limit the influence of large errors. By integrating these loss components, the total temporal loss function is defined as:
$\mathcal{L}_{\text{temporal}}=\alpha\cdot\mathcal{L}_{\text{class}}+\beta\cdot\mathcal{L}_{\text{dice}}+\gamma\cdot\mathcal{L}_{\text{sim}}+\delta\cdot\mathcal{L}_{\text{boundary}}$ (8)
Where $\alpha$, $\beta$, $\gamma$, and $\delta$ are weighting factors that balance the contributions of each loss term. This comprehensive loss function enables the model to effectively capture class imbalances, segment overlaps, temporal dependencies, and boundary precision, leading to improved performance in temporal action segmentation tasks.
4 Experiments and Analysis
Table 2: Class Duration in SAR-RARP and Cholec80 Datasets
Dataset ID Classes/Phases Mean±$\pm$±Std (s)
SAR-RARP 0 Background Classes 12.4 $\pm$ 17.2
1 Picking-up the needle 3.84 $\pm$ 2.66
2 Positioning the needle tip 6.83 $\pm$ 5.65
3 Pushing needle through tissue 7.51 $\pm$ 3.79
4 Pulling the needle out of tissue 6.77 $\pm$ 3.73
5 Tying a knot 26 $\pm$ 7.65
6 Cutting the suture 6.89 $\pm$ 5.31
7 Returning the needle 6.91 $\pm$ 5.41
Cholec80 P1 Preparation 125 $\pm$ 95
P2 Calot triangle dissection 954 $\pm$ 538
P3 Clipping and cutting 168 $\pm$ 152
P4 Gallbladder dissection 857 $\pm$ 551
P5 Gallbladder packaging 98 $\pm$ 53
P6 Cleaning and coagulation 178 $\pm$ 166
P7 Gallbladder retraction 83 $\pm$ 56
Table 3: Comparison of Models Across Datasets (SAR-RARP, Jigsaw, Cholec80)
Dataset Settings Acc Prec Rec F1@10 F1@25 F1@50 Edit SAR-RARP CSAtt 0.61 0.68 0.59 0.56 0.52 0.28 0.73 CHAtt 0.73 0.77 0.71 0.58 0.54 0.40 0.75 CSFWNO 0.82 0.89 0.80 0.77 0.63 0.49 0.74 CSSWSO 0.84 0.87 0.82 0.79 0.65 0.51 0.76 CDWSO 0.87 0.85 0.81 0.85 0.68 0.55 0.79 CHFW 0.84 0.87 0.81 0.80 0.66 0.52 0.76 CHDSWO 0.88 0.89 0.80 0.84 0.72 0.58 0.80 ACSAtt 0.65 0.67 0.61 0.57 0.57 0.43 0.74 ACHAtt 0.68 0.65 0.60 0.64 0.60 0.47 0.80 ACSFWNO 0.87 0.89 0.80 0.83 0.69 0.56 0.79 ACDWSO 0.88 0.91 0.84 0.84 0.71 0.58 0.81 ACHFW 0.88 0.90 0.87 0.82 0.71 0.59 0.81 ACHSSWO 0.89 0.92 0.88 0.85 0.72 0.60 0.82 ACHDSWO 0.90 0.86 0.87 0.87 0.75 0.64 0.84 Cholec80 CSAtt 0.631 0.712 0.621 0.601 0.585 0.472 0.732 CHAtt 0.748 0.799 0.728 0.718 0.604 0.503 0.751 CSFWNO 0.84 0.821 0.83 0.81 0.695 0.583 0.762 CSSWSO 0.856 0.837 0.846 0.826 0.713 0.604 0.78 CDWSO 0.89 0.87 0.86 0.86 0.77 0.64 0.84 CHFW 0.864 0.845 0.854 0.834 0.722 0.614 0.789 CHDSWO 0.878 0.859 0.868 0.848 0.738 0.631 0.805 ACSAtt 0.686 0.767 0.676 0.656 0.547 0.542 0.741 ACHAtt 0.694 0.754 0.684 0.764 0.656 0.553 0.773 ACSFWNO 0.89 0.871 0.88 0.86 0.752 0.648 0.819 ACDWSO 0.902 0.883 0.892 0.872 0.765 0.664 0.832 ACSSWSO 0.91 0.891 0.9 0.88 0.774 0.675 0.841 ACHFW 0.906 0.887 0.896 0.876 0.77 0.67 0.837 ACHDSWO 0.918 0.899 0.908 0.888 0.783 0.686 0.85 Jigsaw CSAtt 0.742 0.823 0.732 0.712 0.598 0.488 0.705 CHAtt 0.759 0.84 0.79 0.729 0.617 0.509 0.748 CSFWNO 0.851 0.832 0.841 0.821 0.708 0.599 0.775 CSSWSO 0.867 0.848 0.857 0.837 0.726 0.620 0.793 CDWSO 0.89 0.88 0.887 0.86 0.76 0.638 0.809 CHFW 0.875 0.856 0.865 0.845 0.735 0.630 0.802 CHDSWO 0.889 0.87 0.879 0.859 0.751 0.647 0.818 ACSAtt 0.797 0.867 0.687 0.867 0.76 0.658 0.827 ACHAtt 0.905 0.886 0.895 0.875 0.769 0.669 0.836 ACSFWNO 0.901 0.882 0.891 0.871 0.765 0.664 0.832 ACDWSO 0.913 0.894 0.903 0.883 0.778 0.680 0.845 ACHFW 0.917 0.898 0.907 0.887 0.783 0.686 0.85 ACHFSWO 0.939 0.94 0.929 0.919 0.796 0.723 0.893 ACHDSWO 0.921 0.902 0.911 0.891 0.787 0.691 0.854
Table 4: Comparison with state-of-the-art models Across Datasets (RARP, Cholec80, JIGSAW)
Dataset Model Acc Edit F1@10 F1@25 F1@50 RARP MS-TCN Li et al. (2020) 69.0 - 70.7 ActionCLIP Psychogyios et al. (2023) 80.4 80.6 MA-TCN Van Amsterdam et al. (2022) 83.4 81.6 81.7 ASRF Ishikawa et al. (2021) $87.5\pm 3.7$ $83.6\pm 1.9$ $80.1\pm 3.2$ $70.8\pm 6.2$ $59.6\pm 3.2$ NETE Yi et al. (2022) $87.8\pm 3.6$ $83.9\pm 1.5$ $78.7\pm 3.6$ $68.5\pm 3.2$ $54.6\pm 2.4$ Asformer Yi et al. (2021) $\boldsymbol{88.8\pm 5.4}$ $\boldsymbol{84.1\pm 3.1}$ $\boldsymbol{84.3\pm 5.3}$ $68.1\pm 3.4$ $59.4\pm 3.1$ MSBATN(Ours) $88.7\pm 3.4$ $84.3\pm 2.7$ $84.0\pm 3.7$ $\boldsymbol{72.4\pm 2.6}$ $\boldsymbol{63.8\pm 1.5}$ Cholec80 ResNet Yi et al. (2022) 78.3 57.8 52.2 PhaseLSTM Neil et al. (2016) 80.7 61.2 62.4 EndoNet Twinanda et al. (2016) 81.9 TeCNO Czempiel et al. (2020) 88.6 ASRF Ishikawa et al. (2021) $82.5\pm 4.5$ $78.5\pm 3.5$ $84.3\pm 3.4$ $70.6\pm 5.7$ $60.1\pm 3.4$ Asformer Yi et al. (2021) $82.8\pm 4.7$ $74.5\pm 6.7$ $84.1\pm 2.4$ $70.8\pm 6.2$ $57.1\pm 3.2$ NETE Yi et al. (2022) $\boldsymbol{92.8\pm 5.0}$ $\boldsymbol{78.7\pm 9.4}$ $82.4\pm 3.4$ $70.1\pm 3.2$ $53.1\pm 4.8$ MSBATN(Ours) $89.4\pm 2.1$ $77.5\pm 8.0$ $\boldsymbol{84.4\pm 4.4}$ $\boldsymbol{73.1\pm 5.2}$ $\boldsymbol{60.2\pm 8.4}$ JIGSAW K-TCN Van Amsterdam et al. (2022) 83.8 86.3 90.4 MA-TCN Van Amsterdam et al. (2022) 86.8 90.1 93.6 ASRF Ishikawa et al. (2021) $87.8\pm 4.5$ $89.0\pm 2.5$ $80.4\pm 6.5$ $70.1\pm 3.6$ $69.8\pm 2.5$ NETE Yi et al. (2022) $90.8\pm 2.1$ $86.5\pm 3.5$ $74.4\pm 7.8$ $65.1\pm 3.7$ $58.6\pm 2.9$ Asformer Yi et al. (2021) $\boldsymbol{92.8\pm 3.3}$ $\boldsymbol{89.8\pm 6.7}$ $83.5\pm 4.2$ $70.1\pm 3.5$ $65.4\pm 1.8$ MSBATN(Ours) $86.1\pm 6.3$ $89.5\pm 6.5$ $\boldsymbol{84.4\pm 5.5}$ $\boldsymbol{76.1\pm 3.1}$ $\boldsymbol{70.7\pm 2.8}$
Table 5: Performance comparison between Start-End Gaussian cosine and BCE Loss
Settings Start-End Gaussian Cosine Loss Binary Cross Entropy Loss
F1@10 F1@25 F1@50 Edit F1@10 F1@25 F1@50 Edit (Bin)
ACDWSO 0.85 0.71 0.53 0.81 0.76 0.6 0.58 0.78
ACHFW 0.83 0.7 0.55 0.78 0.7 0.51 0.45 0.72
ACSSWSO 0.86 0.72 0.6 0.82 0.73 0.62 0.55 0.79
ACHDSWO 0.87 0.74 0.63 0.84 0.77 0.64 0.53 0.8
We extracted every frame (256x 256x3) from the video sequences and used the Dinov2-Reg feature extractor to transform these frames into a 2048-dimensional feature space. These features serve as inputs to our encoder-decoder architecture. The encoder ingests inputs of shape (T, 2048) and processes them through a dilated acausal TCN, effectively reducing the dimensionality to (T,256) while capturing temporal dependencies over varying scales. These embeddings are further refined through ten layers of multi-head attention mechanisms (with 8 heads), incorporating advanced techniques such as sliding window attention for local dependencies and hierarchical attention for multi-scale temporal modelling. This combination enables the encoder to learn both fine-grained and long-range temporal patterns. The encoder produces frame-level predictions for action classes and boundary classes. To refine these initial predictions, three sequential decoders Zhang et al. (2023); Yi et al. (2021); Ishikawa et al. (2021); Yi et al. (2022) are employed. Each decoder shares a similar structure to the encoder but focuses on progressively refining predictions with higher accuracy and temporal smoothness. For optimization, we utilized the Adam optimizer along with the softmax activation function to ensure stable training. In contrast to some existing approaches Yi et al. (2021), positional encoding was omitted, as the hierarchical and sliding attention mechanisms sufficiently capture temporal ordering and dependencies. Our boundary-aware loss function is designed to enhance the model’s segmentation and transition detection capabilities. It combines a weighted Dice loss, which penalizes the model based on Intersection over Union (IoU), with a Gaussian Cosine Similarity loss that emphasizes smooth transitions by modelling centre-to-boundary similarity. This composite loss function ensures robust learning of temporal dynamics and accurate identification of action boundaries. We split the dataset into 80-20 ratios for training testing for Jigsaw and Cholec80. We used dropout and early stopping to handle over-fitting and set the max epoch to 120. We have considered three different datasets such as SAR-RARP50Psychogyios et al. (2023), JIGSAWSGao et al. (2014), Cholec80 Twinanda et al. (2016), and for more details, please refer to supplementary material.
4.1 Evaluation metrics
To evaluate the performance of our models, we used frame-wise accuracy along with segmental-wise overlap. The frame-wise accuracy (FWA) is calculated as FWA=correct framestotal $\text{FWA}=\frac{\text{correct frames}}{\text{total frames}}$ and to asses the temporal performance, we measured the segmental scores, which include the segmental edit distance and the segmental F1 score. The Segmental F1 Score measures the temporal overlap between predicted and ground truth action segments. It is computed as the harmonic mean of precision and recall, focusing on how well the predicted segments align with the actual segments in terms of duration and timing. The segmental F1 score was assessed at overlapping thresholds of 10%, 25%, and 50%, referred to as F1@10, F1@25, and F1@50, respectively. Determine True Positives (TP), False Positives (FP), and False Negatives (FN). A predicted segment is considered a TP if its IoU with a ground truth segment exceeds a predefined threshold $T$. For F1@10%, $T=0.10$, For F1@25%, $T=0.25$, For F1@50%, $T=0.50$. FPs are predicted segments without a matching ground truth segment above the threshold. FNs are ground truth segments without a matching predicted segment above the threshold. Calculate Precision and Recall:
$\text{Precision}=\frac{\text{TP}_{IoU>T}}{\text{TP}_{IoU>T}+\text{FP}_{IoU>T}}$
$\text{Recall}=\frac{\text{TP}_{IoU>T}}{\text{TP}_{IoU>T}+\text{FN}_{IoU>T}}$
Compute Segmental F1 Score:
[…]
The Edit Score assesses the sequence of predicted action segments by comparing it to the ground truth sequence. It is calculated based on the Levenshtein distance, which quantifies the minimum number of edit operations (insertions, deletions, substitutions) required to transform the predicted sequence into the ground truth sequence. A higher Edit Score indicates a closer match between the predicted and actual sequences, reflecting the model’s proficiency in capturing the correct order and transitions of actions. This metric is less sensitive to minor temporal shifts but penalizes over-segmentation and out-of-order predictions. Calculate the Levenshtein distance $D$ between the predicted and ground truth sequences. The Edit Score is then computed as Edit $\text{Edit Score}=1-\frac{D}{\max(L_{p},L_{g})}$ where $L_{p}$ and $L_{g}$ are the lengths of the predicted and ground truth sequences, respectively. The Segmental F1 Score is sensitive to over-segmentation and under-segmentation errors, penalizing models that divide actions into too many segments or fewer segments. In contrast, the Edit Score penalizes incorrect action order and over-segmentation but is more tolerant of slight temporal misalignments. The Segmental F1 Score evaluates the spatial and temporal precision of action boundaries, whereas the Edit Score assesses the structural and sequential accuracy of the predicted action sequence.
4.2 Results Analysis
Our Multi-staged Baundary aware network has three main differences form Asformer, First we employed acausal convolution layer in the initial layers of encoders and decoders. Secondly, we used hierarchical sparse multi-head window attention. Finally, we calculated centre sampled Boundary Aware Loss leverages the strengths of both cosine similarity and Gaussian functions, making it effective for segmentation tasks. We conducted the following experiments to demonstrate the effects of Causal and Acausal convolution along with hierarchical dual or single sliding window attentions.
4.2.1 Effect of Convolution and sliding window attention
We assessed the effectiveness of dilated causal and acausal convolutions, along with various attention settings during training. As illustrated in Table 3, the causal settings include: - Causal Self-Attention (CSAtt) - Causal Hierarchical Attention (CHAtt) - Causal Sliding Fixed Window with No Overlap (CSFWNO) - Causal Sliding Single Window with Overlap (CSSWO) - Causal Dual Sliding Window with Overlap (CDWSO) - Causal Hierarchical Fixed Window with Overlap (CHFWO) - Causal Hierarchical Dual Sliding Window with Overlap (CHDSWO)
In addition, we incorporated acausal convolutions, which include: - Acausal Self-Attention (ACSAtt) - Acausal Hierarchical Attention (ACHAtt) - Acausal Sliding Fixed Window with No Overlap (ACSFWNO) - Acausal Sliding Single Window with Overlap (ACSSWO) - Acausal Dual Sliding Window with Overlap (ACDWSO) - Acausal Hierarchical Fixed Window with Overlap (ACHFWO) - Acausal Hierarchical Dual Sliding Window with Overlap (ACHDSWO) Through ongoing evaluation of these settings, we concluded that for longer sequences, Dual Sliding Overlapping Window Attention outperformed Single Sliding Overlapping Window. Conversely, for shorter action sequences in the Jigsaw dataset, the Single Sliding Window demonstrated superior performance.
Table 3 highlights the effectiveness of proposed hierarchical sliding window attention with causal and acausal convolution. Settings with acausal convolution performed better than causal convolution as causal convolution restricts information flow to the present and past, ensuring that no future data influences the current output. In contrast, acausal convolution permits the output to depend on past, present, and future input values, which is useful during training. We also observe that self-attention and fixed hierarchical attention performed poorly. Vanilla self-attention computes attention scores between all token pairs, leading to a quadratic complexity. Such complexity becomes prohibitive for long sequences, limiting scalability. Hierarchical attention processes input sequences at multiple hierarchical levels with predetermined structures, which may not adapt well to varying local contexts within the data. Hierarchical Sliding Window Sparse Attention focuses on local neighbourhoods dynamically, capturing essential local patterns and structures. Stacking layers of such windowed attention results in a large receptive field, where top layers have access to all input locations and can build representations that incorporate information across the entire input. Single-Sliding Window Attention performs well on the Jigsaw dataset used in surgical training, where samples are shorter and backgrounds are simpler than in real-world datasets like SAR-RARP and Cholec80. In contrast, Dual-Sliding Window Attention performs better in SAR-RARP and Cholec80, which feature complex and dynamic actions. Single-Sliding Window Attention captures short-range dependencies effectively but may struggle with long-range connections, leading to a fragmented understanding in more complex videos. Meanwhile, Dual-Sliding Window Attention is better at capturing both short- and long-range dependencies, making it suitable for handling diverse action durations and transitions.
4.2.2 Effect of Start-End Boundary Aware Loss Function
Traditional loss functions, such as Binary Cross-Entropy (BCE), predominantly focus on frame-wise classification, which can result in either over-segmentation or under-segmentation. This approach has inherent limitations, particularly when addressing complex or ambiguous boundaries. Additionally, binary boundary classification may lead to instability during training due to substantial variations in loss values both within and between classes. The results of the proposed boundary-aware loss function show that our proposed approach effectively handled these challenges by capitalizing on the distribution of similarity scores between frames generated by decoders. Table 5 demonstrates the effectiveness of our method, which incorporates actual convolution and sliding window attention settings, delivering improved results across the five metrics outlined in the previous section compared to the binary-boundary-based approach.
5 Conclusion and Feature Work
Segmenting actions in lengthy surgical settings is challenging due to the complex interactions between tools and organs. There are no established standards for optimal performance, as most assessments rely on frame-wise recognition, which has now shifted to segment-wise understanding. In this article, we present a TCN-transformer-based framework with hierarchical sliding window attention that operates with linear computational costs.
We found that selecting the right attention module can be difficult, and neighbouring frames often lead to ambiguous results, causing over- and under-segmentation in frame-wise segmentation. Segment-wise refinement is more effective in improving understanding of surgical activities. Additionally, binary label boundaries are often inadequate because the start, centre, and end of actions vary by individual expertise, leading to poor performance. Our proposed framework addresses these issues with hierarchical sliding window attention for global-local context and includes a novel Boundary Segmentation Loss to better recognize central and boundary points. Our boundary-aware segmentation approach outperforms traditional binary detection methods. Experimental results show that our model matches or exceeds the performance of current state-of-the-art methods across various metrics, ensuring robustness in complex surgical environments.
These results open new interesting research directions, in particular, exploring anticipation tasks where decoders predict the next segment while accounting for uncertainty in a goal-driven manner. Furthermore, integrating instructional prompts to enhance and refine the learning process represents another promising avenue for future investigation.

### v2

@@
-Conceptualization of this study, Methodology, Writing
+
 Abstract
-Understanding actions within surgical workflows is essential for evaluating post-operative outcomes.
-However, capturing long sequences of actions performed in surgical settings poses challenges, as individual surgeons have their unique approaches shaped by their expertise, leading to significant variability.
-To tackle this complex problem, we focused on segmentation with precise boundaries, a demanding task due to the inherent variability in action durations and the subtle transitions often observed in untrimmed videos.
-These transitions, marked by ambiguous starting and ending points, complicate the segmentation process.
-Traditional models, such as MS-TCN, which depend on large receptive fields, frequently face challenges of over-segmentation (resulting in fragmented segments) or under-segmentation (merging distinct actions).
-Both of these issues negatively impact the quality of segmentation.
-To overcome these challenges, we present the Multi-Stage Boundary-Aware Transformer Network (MSBATN) with hierarchical sliding window attention, designed to enhance action segmentation.
-Our proposed approach incorporates a novel unified loss function that treats action classification and boundary detection as distinct yet interdependent tasks.
-Unlike traditional binary boundary detection methods, our boundary voting mechanism accurately identifies start and end points by leveraging contextual information.
-Extensive experiments using three challenging surgical datasets demonstrate the superior performance of the proposed method, achieving state-of-the-art results in F1 scores at thresholds of 25% and 50%, while also delivering comparable performance in other metrics.
+Understanding actions within surgical workflows is critical for evaluating post-operative outcomes and enhancing surgical training and efficiency.
+Capturing and analyzing long sequences of actions in surgical settings is challenging due to the inherent variability in individual surgeon approaches, which are shaped by their expertise and preferences.
+This variability complicates the identification and segmentation of distinct actions with ambiguous boundary start and end points.
+The traditional models, such as MS-TCN, which rely on large receptive fields, that causes over-segmentation, or under-segmentation, where distinct actions are incorrectly aligned.
+To address these challenges, we propose the Multi-Stage Boundary-Aware Transformer Network (MSBATN) with hierarchical sliding window attention to improve action segmentation.
+Our approach effectively manages the complexity of varying action durations and subtle transitions by accurately identifying start and end action boundaries in untrimmed surgical videos.
+MSBATN introduces a novel unified loss function that optimises action classification and boundary detection as interconnected tasks.
+Unlike conventional binary boundary detection methods, our innovative boundary weighing mechanism leverages contextual information to precisely identify action boundaries.
+Extensive experiments on three challenging surgical datasets demonstrate that MSBATN achieves state-of-the-art performance, with superior F1 scores at 25% and 50% thresholds and competitive results across other metrics.
 keywords:
@@
 1 Introduction
-Surgical action segmentation involves analyzing complex elements such as instruments, organs, fluids, or smoke to identify and categorize distinct actions or phases in surgical videos Mascagni et al.
+Surgical action segmentation involves analyzing instruments, organs, fluids, and smoke to identify and categorise distinct actions or phases in complex surgical videos Mascagni et al.
 (2022).
-This task is crucial for enhancing surgical training by providing detailed feedback to trainees, improving workflow efficiency and patient safety through real-time monitoring and optimization of procedures, and advancing the performance of autonomous or semi-autonomous surgical robots capable of understanding and executing complex tasks.
-Recent advancements in this domain have led to the creation of sophisticated architectures tailored to the unique challenges of surgical environments (like Gazebo, ROS2 software development kit (SDK), etc), enabling better decision-making services during operations and driving progress in surgical robotics and education Kiyasseh et al.
+This process is essential to improve surgical training by providing detailed feedback to trainees and advancing the performance of autonomous surgical robots.
+These robots can understand and perform intricate tasks, thus improving workflow efficiency and patient safety through real-time monitoring and optimisation of procedures.
+Recent advances in this field have led to the development of sophisticated architectures tailored to the unique challenges of surgical environments, such as Gazebo and the ROS2 software development kit (SDK).
+These innovations support better decision-making during operations and drive progress in surgical robotics and education Kiyasseh et al.
 (2023).
-Surgical videos are often heavily annotated manually, where each frame is labelled to a surgical phase or sub-phase or specific action/gesture, as well as the surgical instruments present in that frame Psychogyios et al.
+Figure 1: The action-transition graph for SAR-RARP50 features nodes 1 to 7, representing canonical suturing actions.
+A typical sequence is: Pick up the needle (1) → Position the tip (2) → Push through the tissue (3) → Pull out (4).
+After this sequence, the surgeon can either return or drop the needle (7), tie a knot (5), cut the suture (6), or skip the last two steps.
+This graph allows for loops and re-entries, so it doesn’t follow a single linear path from start to finish.
+Surgical videos are typically annotated manually, with each frame labeled according to the surgical phase, sub-phase, or specific actions/gestures, as well as the instruments present in that frame Psychogyios et al.
 (2023); Twinanda et al.
 (2016).
-Recent datasets have incorporated action triplets Nwoye et al.
-(2022), which include action, instrument, and organ information for frame-wise action recognition and instruments or organs segmentation Nwoye et al.
+Recent datasets have introduced action triplets Nwoye et al.
+(2022), which include information about the action, instruments, and organs for frame-wise action recognition and segmentation of instruments or organs Nwoye et al.
 (2023); Chadebecq et al.
 (2023).
-However, temporal evolution hierarchies are not well defined as they are subjective in nature.
-The high to low-level temporal component of a surgical operation or intraoperative activity is termed phase, steps and tasks.
-For more details please refer attached supplementary material file.
-Figure 1 demonstrates an effective approach to handle temporal variations by modelling start-end transition in long videos.
-It starts with frame-wise action and boundary recognition, and then it goes through a segment-wise robust refinement process to avoid over-under segmentation.
-For efficient analysis of surgical workflows and improving postoperative outcomes, a robust understanding of both spatial context and temporal relationships, the arrangement and transitions between actions or gestures, are essential Chadebecq et al.
+Most of these datasets focus primarily on spatial information, while providing minimal temporal information.
+There is a standard temporal hierarchy that surgeons follow during procedures.
+The highest-level temporal component of a surgical operation or intraoperative activity is referred to as a "phase," which encompasses a series of related activities.
+Each phase consists of several "steps," which are the specific actions taken to achieve the goals of that phase.
+Each step represents a procedure-specific segment aimed at addressing the patient’s condition, while the "task" is a sub-component of a step that involves a series of actions Meireles et al.
 (2021).
-In this regard, most publicly available datasets often lack these temporal hierarchies and their evolutions.
-Additionally, similarities between different phases, including unclear or ambiguous starting, ending and transitions between actions 2, lead to inaccurate recognition.
-Figure 1: Overview of the proposed framework, which focuses on segments encompassing the action’s start, centre, and end, enabling the model to capture nuanced transitions and improve boundary precision.
-For improved surgical video analysis, most existing models are built based on deep neural networks which are broadly divided into single-stage and multi-stage where the architectures evolved from RNN to Temporal Convolution Network(TCN)Farha and Gall (2019); Li et al.
+However, the definitions of temporal evolution hierarchies can be subjective and depend on the surgeon’s experience, which is why they are rarely disclosed alongside the videos.
+This lack of structure, coupled with visual similarities between actions and fuzzy boundaries, often leads to ambiguous transitions (Figure 1).
+For example, in SAR-RARP50, the action-transition graph is highly cyclic: any of the seven suturing gestures can be repeated or skipped, resulting in frequent re-entries and out-of-order sequences.
+Deep learning pipelines for surgical video analysis have advanced significantly, evolving from recurrent networks to temporal convolutions, and more recently, to Transformers Farha and Gall (2019); Li et al.
 (2020); Czempiel et al.
 (2020); Yi et al.
-(2022), and Transformer Yang et al.
+(2022); Yang et al.
 (2024); Zhang et al.
@@
 (2022); Shi et al.
-(2023) based approaches.
-Single-stage models make predictions directly from visual-temporal features.
-These models utilize a small window of time around the target frame to capture spatial-temporal features, which minimizes computational costs.
-Early models like EndoNetTwinanda et al.
-(2016); Twinanda (2017) focused on detecting tool presence and recognizing phases using RNN-based methods.
-EndoNetTwinanda et al.
-(2016) is the first method proposed for two tasks simultaneously in a multitasking manner, including tool presence detection and phase recognition.
-EndoLSTM Twinanda (2017), as the improved EndoNet, replaced hierarchical hidden Markov models (HHMMs) with LSTM for temporal refinement and improved the results substantially.
-Besides, in Mondal et al.
-(2019), a multitask deep learning framework was proposed comprising ResNet-50 and Bi-LSTM.
-It performs better in tool detection and phase recognition.
-In Jin et al.
-(2020), a single-staged multitask recurrent convolutional network was proposed with correlation loss (MTRCNet-CL) for surgical video analysis.
-This multitask learning method used a new correlation loss to minimize tool and phase prediction divergences.
-However, RNN and LSTM-based approaches often struggle to capture long-term dependencies that are essential for accurately modelling the entire surgical process.
-For instance, models like Actionformer Zhang et al.
+(2023).
+Early systems, such as EndoNet Twinanda et al.
+(2016) and EndoLSTM Twinanda (2017), utilised CNN backbones for frame features and RNN modules for temporal smoothing, achieving the first joint modeling of tool presence and phase recognition.
+Subsequent variants incorporated correlation losses, bidirectional LSTMs, or multitask heads to improve accuracy Mondal et al.
+(2019); Jin et al.
+(2020).
+However, RNN-based designs often struggle to retain dependencies over minutes—to hours—long procedure.
+To expand the temporal receptive field without encountering the vanishing gradient problems associated with RNNs, researchers have shifted to Temporal Convolutional Networks (TCNs)) Farha and Gall (2019); Li et al.
+(2020); Czempiel et al.
+(2020); Yi et al.
+(2022).
+The MS-TCN family employs dilated 1-D kernels that efficiently aggregate historical information while maintaining a low inference latency Farha and Gall (2019); Li et al.
+(2020).
+Recently, single-stage Transformers like Actionformer Zhang et al.
 (2022) and TriDet Shi et al.
-(2023) integrate Transformer architecture for action recognition.
-However, these single-stage models are designed for small actions while each phase or action segment in surgery takes several minutes, or the whole procedure sometimes takes more than 1 hour.
-Because of this research, the signal-stage models are not ideal for long untrimmed video analysis.
-Multi-stage models, on the other hand, include an additional refinement stage to improve the initial predictions by correcting the errors.
-Multi-stage models are considered more suitable for surgical phase recognition to optimize the complexity and ordered nature of surgical activitiesCzempiel et al.
-(2020); Van Amsterdam et al.
+(2023) have provided global self-attention over short clips, excelling in fine-grained gesture localisation.
+However, these single-pass models were designed for brief human actions; their fixed window makes them unsuitable to capture the hierarchical nature of surgical workflows.
+Multi-stage architectures address this limitation by incorporating one or more refinement passes to correct over- or under-segmentation produced in the first stage.
+TeCNO Czempiel et al.
+(2020) chains several TCN blocks, each trained to refine the output of its predecessor, while MTMS-TCN jointly learns phases and steps to leverage shared cues Farha and Gall (2019); Li et al.
+(2020); Ramesh et al.
 (2021).
-Multi-stage models such as TeCNOCzempiel et al.
-(2020) and Farha and Gall (2019); Li et al.
-(2020) employ multiple temporal convolutional networks to refine predictions in subsequent stages.
-However, these methods often neglect high-level semantic information and struggle to refine predictions effectively in an end-to-end manner, leading to over or under-segmentation errors.
-This is because the refinement stage focuses on adjusting initial predictions without fully understanding the actions involved during training, and errors remain during inference.
-This can lead to misinterpretations in complex scenarios where context is crucial.
-Recently, in Ramesh et al.
-(2021), a multitask multistage temporal convolutional network (MTMS-TCN) was proposed to jointly learn the tasks of phase and step recognition.
-Some studies applied sensor data with attention mechanism to enhance the understanding of surgical processes Van Amsterdam et al.
+Other research has integrated kinematic sensor streams or attention mechanisms to improve temporal reasoning Van Amsterdam et al.
 (2022); Goldbraikh et al.
-(2023), however aligning sensor(kinematic) data with corresponding video frames is complex due to differences in sampling rates and potential time lags(different frequencies), leading to misalignment issues.
-To improve the segmentation, some research has focused on incorporating segment-level modeling with transformer Ding and Li (2022); Jin et al.
+(2023), and segment-level Transformers have refined ambiguous boundaries by analysing high-level proposals Ding and Li (2022); Jin et al.
 (2022).
-Their main idea is to use high-level segments from surgical videos to clarify ambiguous semantics presented in neighboring frames.
-Figure 2: Ambiguous action ordering in untrimmed long surgical videos
-To overcome the above limitations, the current methods use sliding window detectors Czempiel et al.
-(2020), but they struggle to capture long-term patterns effectively Twinanda et al.
-(2016); Twinanda (2017); Mondal et al.
-(2019); Czempiel et al.
-(2020).
-Additionally, these approaches overlook the finer details that are essential for understanding the intricate relationships between spatial and temporal features Kiyasseh et al.
+Despite these advancements, refinement modules often work with hard predictions without an explicit model of surgical semantics, which can lead to persistent systematic errors, particularly when actions visually resemble one another.
+Figure 2: Multi-stage Boundary-Aware Refinement Pipeline: The initial prediction with typical over- and under-segmentation errors (segments are split too finely, merged, or slightly shifted) misleads action order.
+Our approach resolves this by applying a Gaussian centre-weighting scheme to Action and Boundary segments to minimise the error.
+The refined prediction aligns with the annotated ground truth, and boundaries without re-processing the video.
+Furthermore, these approaches overlook the finer details that are essential for understanding the intricate relationships between spatial and temporal features Kiyasseh et al.
 (2023), resulting in over and under-segmentation errors.
 The over-segmentation occurs when a segment is incorrectly divided into multiple action segments, leading to the detection of more actions than actually exist.
-These approaches are overly sensitive to minor variations in the input data, causing them to interpret slight changes as new action boundaries Ishikawa et al.
+These approaches are overly sensitive to minor variations in the input data, leading them to interpret slight changes as new action boundaries Ishikawa et al.
 (2021).
 On the other hand, under-segmentation happens when distinct actions are mistakenly merged into a single segment.
-This typically arises due to a lack of sufficient temporal resolution or failure to recognize subtle transitions between different actions.
-So, the model overlooks actual action boundaries, resulting in fewer detected action segments than actually exist.
-Most of the approaches fail to model the continuous and smooth transitions characteristic of surgical activities due to treating actions and boundaries as discrete events without accounting for their gradual shifts between actions Fard et al.
+This typically arises due to a lack of sufficient temporal resolution or the inability to recognise subtle transitions between different actions.
+So, the model overlooks the actual action boundaries, resulting in fewer detected action segments than actually exist.
+The continuous and smooth transitions characteristic of surgical activities are often ignored due to treating actions and boundaries as discrete events without accounting for their gradual shifts between actions Fard et al.
 (2017); Ishikawa et al.
@@
 (2021).
-To address these issues, first, we propose a framework that extracts spatial features with DinoV2 by discarding irrelevant large patch tokens with register tokens Oquab et al.
+Our solution (Fig.
+2) tackles these problems in two steps: firstly, a backbone network predicts initial labels for actions and their boundaries.
+Secondly, a lightweight post-processing module corrects over- and under-segmentation by focusing exclusively on the start–end transitions.
+Because it operates only on predicted segments, it adds negligible computational cost to the main model.
+Training utilises a duration-aware loss weighting that adjusts each segment’s contribution inversely based on its duration, ensuring that short yet clinically significant actions are not overshadowed by longer ones.
+We propose a framework that extracts spatial features with DinoV2 by discarding irrelevant large patch tokens with register tokens Oquab et al.
 (2023).
-Second, for temporal modelling, we present an encoder-decoder model that is built based on the combination of dilated TCN, unified hierarchical and sliding window attention, which captures coarse-to-fine context.
-The hierarchical representation with dynamic sliding window attention is used to handle long input sequences that organize the attention layers hierarchically to first focus on local relationships and then gradually capture global dependencies.
-Last, we develop a unified loss function by combining the loss functions of individual tasks into a single action segmentation to achieve effective boundary detection.
-This holistic approach allows one to learn multiple tasks simultaneously, leveraging shared representations to improve overall segmentation performance.
+For temporal modelling, we present an encoder-decoder architecture that is built based on the combination of dilated TCN, unified hierarchical and sliding window attention, which captures fine-to-coarse context.
+The hierarchical representation with dynamic sliding window attention is used to handle long input sequences that organise the attention layers hierarchically to first focus on local relationships and then gradually capture global dependencies.
+We design a unified loss function by combining the loss functions of individual tasks into a single action segmentation to achieve effective boundary detection.
 Our approach introduces the following key innovations:
   • We present a multi-stage TCN-Transformer architecture with dilated acausal convolution to widen the receptive field that predicts and refines both actions and boundaries.
-  • We introduce hierarchical sliding window attention mechanisms (HSWA) to resolve ambiguities among action orders from different temporal resolution.
+  • We introduce hierarchical sliding window attention mechanisms (HSWA) to resolve ambiguities among action orders from different temporal resolutions.
 It links high-level action segment structures with fine-grained frame-wise details, enabling the model to better understand long-term temporal dependencies and immediate frame-level changes.
-  • We propose a boundary aware loss function that gradually measures cosine similarity from center to boundaries inside action segments that assist in minimizing over or under-segmentation error.
+  • We propose a boundary-aware loss function that gradually measures cosine similarity from centre to boundaries inside action segments that assist in minimising over- or under-segmentation error.
 The remainder of this article is structured as follows: Section 2 focuses on related work.
@@
 (2016) was developed for phase recognition in cholecystectomy videos by learning inherent visual features through a CNN architecture.
-Recognizing the importance of tool information in phase recognition, EndoNet adopted a multi-task framework that incorporates tool presence detection into the feature learning process.
-An extension of AlexNet was designed with dual objectives: one part focused on detecting the presence of surgical tools, while the other was responsible for recognizing seven distinct surgical phases.
+Recognising the importance of tool information in phase recognition, EndoNet adopted a multi-task framework that incorporates tool presence detection into the feature learning process.
+An extension of AlexNet was designed with dual objectives: one part focused on detecting the presence of surgical tools, while the other was responsible for recognising seven distinct surgical phases.
 To capture temporal dependencies, EndoNet integrated a Hierarchical Hidden Markov Model (HMM), enabling the modelling of temporal transitions between phases.
@@
 (2019); Yi et al.
-(2021), which accommodates variable-length inputs and outputs, making it well-suited for detecting overlapping or sequential surgical tasks.
+(2021), which accommodates variable-length inputs and outputs, making it well-suited for detecting overlapping surgical tasks.
 Temporal Action Detection (TAD) methodologies are mainly classified into two-stage and one-stage-based methods.
@@
 (2022).
-Although this method often achieves higher accuracy due to its detailed localization process, it tends to be computationally intensive, posing challenges for real-time applications Yi et al.
+Although this method often achieves higher accuracy due to its detailed localisation process, it tends to be computationally intensive, posing challenges for real-time applications Yi et al.
 (2022).
-On the other hand, the one-stage approach together consolidates action proposal generation and classification in a single framework, which offers computational efficiency for real-time scenarios.
-For instance, models like ActionFormer Zhang et al.
+On the other hand, the one-stage approach consolidates action proposal generation and classification in a single framework, which offers computational efficiency for real-time scenarios.
+Models like ActionFormer Zhang et al.
 (2022), and Surgformer Yang et al.
-(2024) are transformer-based architectures that have achieved enhanced precision of action localization while maintaining computational efficiency.
+(2024) are transformer-based architectures that have achieved enhanced precision of action localisation while maintaining computational efficiency.
 These models capture long-range dependencies and contextual information that helps improve the accuracy of action segmentation and detection.
-However, one-stage methods have historically faced difficulties in precisely localizing action boundaries in long untrimmed videos due to their small receptive fields, which is crucial for optimal performance in TAD.
+However, one-stage methods have historically faced difficulties in precisely localising action boundaries in long untrimmed videos due to their small receptive fields, which is crucial for optimal performance in TAD.
 2.2 Boundary Detection
@@
 VB inference approximated the full posterior distribution, which improved model flexibility and reduced redundant components.
-Label assignment and tracking were achieved through Kullback-Leibler (KL) distance, ensuring consistent tracking across video clips.
-In contrast to rigid boundary detection methods, the Soft-Boundary Unsupervised Gesture Segmentation (Soft-UGS) Fard et al.
-(2017) approach was proposed as a soft-boundary clustering technique that allowed for gradual transitions between actions.
-This method utilized kinematic trajectory data, Dynamic Time Warping (DTW), and probabilistic clustering to segment surgical gestures effectively in real-time applications.
-While these probabilistic models laid the groundwork for more advanced methods, they had limitations in precision, particularly in detecting abrupt transitions in video content.
-On the other hand, the advent of deep learning introduced new techniques for addressing the temporal and spatial complexities of action boundary detection.
-For the surgical domain, sequence models were applied to map variable-length input sequences to output sequences while incorporating attention mechanisms for more accurate boundary localization.
+Label assignment and tracking were accomplished using Kullback-Leibler (KL) distance, ensuring consistent tracking across video clips.
+Unlike rigid boundary detection methods, the Soft-Boundary Unsupervised Gesture Segmentation (Soft-UGS) approach Fard et al.
+(2017) was proposed as a soft-boundary clustering technique that allowed for gradual transitions between actions.
+This method effectively utilised kinematic trajectory data, Dynamic Time Warping (DTW), and probabilistic clustering to segment surgical gestures in real-time applications.
+Although these probabilistic models provided a foundation for more advanced techniques, they faced limitations in precision, especially in detecting abrupt transitions in video content.
+In contrast, the emergence of deep learning introduced new techniques for addressing the temporal and spatial complexities of action boundary detection.
+For the surgical domain, sequence models were applied to map variable-length input sequences to output sequences while incorporating attention mechanisms for more accurate boundary localisation.
 A double attention mechanism(such as Namazi et al.
@@
 ASRF refines action segmentation through two branches: one for action predictions and another for boundary regression.
-By focusing specifically on predicted action boundaries, ASRF achieved overall action localization accuracy.
+By focusing specifically on predicted action boundaries, ASRF achieved overall action localisation accuracy.
 ASRF* Yi et al.
@@
 TriDet introduced the Trident-head detection mechanism, which models action boundaries by estimating relative probability distributions instead of directly predicting offsets.
-This innovation significantly improved boundary localization accuracy while reducing computational overhead.
+This innovation significantly improved boundary localisation accuracy while reducing computational overhead.
 EMB tackles uncertainties in temporal boundaries by transforming rigid and manually labelled endpoints into an elastic set of boundaries; however, it does tend to generate a high number of false positives.
 Building upon these advancements, our action boundary network employs a voting mechanism to improve segmentation by accurately finding the start and end of an action.
-By utilizing the predicted action boundaries, our framework refines frame-wise predictions, ultimately improving the performance of action segmentation.
-Table 1: Notation table
-Symbol Description
-$T$ Sequence length (number of frames).
-$D$ Dimensionality of frame features.
-$\mathbf{x}_{t}$ Feature vector of the $t$-th frame.
-$C$ Set of action categories (including background).
-$S$ Set of action segments.
-$W$ Base window size for sliding attention.
-$d$ Dilation factor for dilated attention.
-$T_{s}$ Sequence length at scale $s$.
-$\alpha_{ij}^{s}$
-Attention weight for token $x_{i}$ attending
-to $x_{j}$ at scale $T_{s}$.
-$\mathcal{L}_{\text{dice}}$ Dice loss for segment overlap.
-$\mathcal{L}_{\text{sim}}$ Gaussian-weighted cosine similarity loss.
-$\mathcal{L}_{\text{boundary}}$ Boundary-aware loss for segment precision.
-$\alpha,\beta,\gamma,\delta$ Weights for different loss components.
-$G(t,t+1)$ Gaussian weight emphasizing temporal proximity.
-$G_{\text{boundary}}(t)$ Gaussian weight centered on segment boundaries.
+By utilising the predicted action boundaries, our framework refines frame-wise predictions, ultimately improving the performance of action segmentation.
 3 Proposed Framework
@@
 Formally, given a sequence of $T$-frame features $X=[\mathbf{x}_{1},\mathbf{x}_{2},\ldots,\mathbf{x}_{T}]\in\mathbb{R}^{T\times d},\quad\text{where }\mathbf{x}_{T}\in\mathbb{R}^{d}$ represents the $d$-dimensional feature of the $t$-th frame.
-The objective is to predict a sequence of actions $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{T}],\quad\text{where }\hat{y}_{t}\in C.$ The set $C=\{0,1,2,\ldots,C,-100\}$ includes all action categories in the dataset, with $-100$ representing the background class.
-The frame-wise ground truth labels are $Y=[y_{1},y_{2},\ldots,y_{T}],$ accompanied by segment information $S=\{(y_{s1},y_{e1},a_{c1}),\ldots,(y_{sn},y_{en},a_{cn})\},$ where $y_{s}$ and $y_{e}$ denote the start and end of a segment, $a_{c}$ represents the action class label, and $n$ is the number of segments, including background segments and notations are listed in Table 1.
+The objective is to predict a sequence of actions $\hat{Y}=[\hat{y}_{1},\hat{y}_{2},\ldots,\hat{y}_{T}],\quad\text{where }\hat{y}_{T}\in C.$ The set $C=\{0,1,2,\ldots,c\}$ includes all action categories in the dataset.
+The frame-wise ground truth labels are $Y=[y_{1},y_{2},\ldots,y_{T}],$ accompanied by segment information $S=\{(y_{s_{1}},y_{e_{1}},a_{c_{1}}),\ldots,(y_{s_{n}},y_{e_{n}},a_{c_{n}})\},$ where $y_{s}$ and $y_{e}$ denote the start and end of a segment, $a_{c}$ represents the action class label, and $n$ is the number of segments, including background segments.
 The goal is to accurately predict action transitions and segment boundaries.
 Since action transitions are continuous and boundaries are often subjective, considering this 5% of the total segment length is labeled as a transition buffer zone at the start and end of each segment.
-Our proposed framework is depicted in figure 3 that enables three distinct parts.
-First, the frame wise feature extraction, second, Encoder-Decoder architecture with hierarchical sliding window attention, and last, boundary aware loss function that guides to detect the transition with start-center-end points.
-Figure 3: Proposed Multi-Stage Boundary-Aware Transformer Network (MSBATN) architecture
+Our proposed framework, the Multi-Stage Boundary-Aware Transformer Network (MSBATN), is illustrated in Figure 3.
+MSBATN is designed to improve action segmentation in untrimmed surgical videos by accurately identifying action boundaries, even in lengthy and complex sequences.
+It achieves this through three main components: robust frame-wise feature extraction, an encoder-decoder architecture based on TCN-Transformer-TCN that incorporates Hierarchical Sliding Window Attention (HSWA) and specialised acausal/causal convolutions, and a novel boundary-aware loss function.
+Figure 3: Proposed Multi-Stage Boundary-Aware Transformer Network (MSBATN): Input video features $(X)$ are first processed by TCN encoder consisting Acausal Conv1D (see right bottom) for temporal feature extraction to produce $X_{tcn}$.
+These features are then fed into the Transformer Encoder, consisting of Hierarchical Sliding-Window Attention (HSWA) (see right top), to generate $X_{encoded}$.
+The $X_{encoded}$ features are expanded to the original temporal resolution to separate Action and Boundary segmentation heads.
+Finally, a TCN Decoder refines these initial predictions to output the final action segments and boundaries.
 3.1 Frame-Wise Feature Extraction
-Dino with Register Oquab et al. (2023) (Dinov2-Reg) captures complex spatial features in a self-supervised manner, Figure 4 depicts the performance in comparison with ViT and Resnet.
+The pipeline begins with an input surgical video from which individual frames are extracted.
+For MSBATN, we specifically use Dinov2-Reg, a feature extractor that improves upon standard ViTs.
+In standard ViTs, images are divided into patches represented by tokens, but "outlier tokens" can introduce errors due to unusually high values.
+Dinov2-Reg addresses this by adding register tokens as dedicated placeholders, enhancing feature robustness and reducing reliance on outlier tokens.
+As shown in Figure 4, Dinov2-Reg Oquab et al.
+(2023) captures complex spatial features more effectively in a self-supervised manner compared to standard ViT and ResNet methods.
 Figure 4: Comparison of frame feature extractors.
-Dinov2-Reg model extracts features using a self-supervised vision transformer model, which processes image-like inputs (frames in this case) to output dense, high-dimensional feature representations.
-The input feature dimensions (256, 1536), where 256 denotes the 16x16 patch dimension, and 1536 is the feature dimension.
-A 2D convolutional layer with a kernel size of (256, 1536) is applied to the input frames, which projects the high-dimensional DINO features (256x1536) into a lower-dimensional representation (1x2048).
-This is equivalent to performing dimensionality reduction via a convolutional operation, which retains important information while reducing the computational complexity for further processing in a Multi-Stage Boundary-Aware Transformer Network (MSBATN).
-For more details, please refer to the attached supplementary material file.
+The Dinov2-Reg extractor converts each frame into a high-dimensional patch-based feature representation (e.g., resulting in a ($H^{\prime}\times W^{\prime}\times C_{dino}$) tensor.
+We use (256, 1536), where 256 denotes the $16\times 16$ patch dimension, and 1536 is the feature dimension.
+To create a fixed-size feature vector per frame suitable for our temporal model, we project these features.
+Specifically, we apply a 2D convolutional layer (or an equivalent linear projection on flattened patch features) to transform the DINOv2 output for each frame into a (d)-dimensional vector (d=2048).
+This results in an initial feature sequence X with dimensions (T, D), where T is the sequence length.
+This step standardises the frame features and prepares them for the subsequent MSBATN architecture.
 3.2 Multi-Stage Boundary-Aware Transformer Network (MSBATN)
-We present MSBATN transformer-based encoder-decoder architecture without positional encoding to improve the performance of the segmentation task.
-Each encoder-decoder block comprises a series of TCN and hierarchical sliding window attention modules.
+The core of our framework is the MSBATN, a Transformer-inspired encoder-decoder architecture designed for effective temporal feature learning and segmentation (see Figure 3).
+It includes an initial TCN-based feature processor, a Transformer-based encoder for deeper contextual understanding, and TCN-based decoders for refinement.
+Notably, MSBATN does not need explicit positional encodings, as its design naturally captures sequence order and relationships through temporal convolutions and attention mechanisms.
 3.2.1 Encoder-Decoder
-The frame-wise features $(T,D)$ are passed through an encoder block.
-The encoder block comprises a series of dilated Acausal Temporal Convolution feed-forward layer blocks to capture context from past and future, followed by the hierarchical sliding window attention module and, lastly, project the encoded features.
-Instead of using traditional fully connected layers, dilated temporal convolutions are employed to expand the receptive field without increasing the number of parameters or sacrificing resolution.
-It introduces local inductive biases to effectively capture temporal dependencies and patterns over extended sequences.
-The model adopts a hierarchical pattern in which the attention layers simultaneously focus on local-global information, mirroring techniques utilized in longformerBeltagy et al.
+The encoder path begins by processing the frame-wise features $X=(B,D_{in},T_{orig})$, where B is batch size, $D_{in}$ is the input feature dimension, and $T$ is the original sequence length.
+The input features first pass through an initial TCN stage.
+It projects the input dimension (2048) to the model’s working dimension (256).
+It employs dilated 1D Acausal TCN where the kernel k is centred on the current frame, covering T-K T, T+K.
+In Causal TCN, the kernel K shifts from right to left, that is, T-K to T thus never peeks into the future, and therefore safe for real-time or streaming inference (Figure 3).
+When the sequence is too long we applied stride greater than 1 to reduce the sequence length $T$ to $T_{reduced}$ resulting in $X_{tcn}=(B,256,T_{reduced})$.
+This makes subsequent attention computations more efficient.
+The dilated convolutions allow the network to build a hierarchical receptive field, capturing local temporal patterns and dependencies over extended sequences effectively.
+These TCN blocks can access both past and local future context within their receptive field during training, enabling robust local feature extraction.
+$X_{tcn}$ is then processed by a stack of encoder blocks.
+Each block aims to capture more complex and longer-range temporal dependencies.
+Each encoder block consists of Dual Sliding Window Attention (DSWA), Hierarchical Temporal Attention (HTA) and Feed-Forward Network (MLP) with Layer Normalisation and Residual Connections.
+Lastly the initial predictions from the encoder are refined by the decoder.
+Instead of acausal convolution in the encoder TCN, the decoder uses causal convolution to avoid information about the future.
+The network is comprised of one encoder and three decoders, with each encoder and decoder containing ten TCN blocks.
+We set the feature dimension in the encoder and decoder layers to 256.
+Additionally, a temporal dropout layer is applied to the input feature of the encoder, which randomly drops the entire feature channel with a dropout rate of 0.3 to mitigate the over-fitting issue.
+3.2.2 Hierarchical Sliding Window Attention
+We present a Hierarchical Sliding Window Attention (HSWA) module that effectively captures temporal dependencies by combining dynamic dual-sliding window attention (DSWA) with hierarchical temporal attention for global context.
+In DSWA, instead of dense self-attention, we use a sparse attention approach inspired by Longformer Beltagy et al.
 (2020).
-To capture the global context, we calculated the hierarchical attention at different scales, and then we aggregated to obtain the global context.
-A detailed discussion is added in the next subsection.
-The decoder considered the temporal relationships between action-boundary segments to improve the initial predictions.
-Similar to the encoder block, the decoder begins with a temporal convolutions block, attention module and output.
-However, instead of acausal convolution, it has causal convolution, so it does not have any information about the future.
-Each attention block uses a weighted residual connection to combine the outputs from the feed-forward layers, with the weight $\alpha$ decreasing exponentially across decoders to prevent error accumulation.
-Last, the decoder generates action boundary labels.
-The network is comprised of one encoder and three decoders, with each encoder and decoder containing ten TCN-attention blocks.
-We set the feature dimension in the encoder and decoder layer to 256.
-Additionally, a temporal dropout layer is applied to the input feature of the encoder, which randomly drops the entire feature channel with a dropout rate of 0.3 to mitigate the over-fitting issue.
-3.2.2 Hierarchical dual Sliding Window Attention
-Instead of computing attention scores across all token pairs, as in self-attention, we employed a sparse-attention mechanism.
-This approach uses a hierarchical sliding window that restricts attention to a fixed-size window around each token at a specific layer.
-The window slides and aggregates scores, meaning each token attends only to its neighboring tokens within the specified window.
-This reduces both space and computational complexity to $O(n)$.
-To capture temporal dependencies at varying distances, a dilated attention window is applied with increasing dilation rates $(0,1,2,3,\dots)$ in each layer.
-This approach allows for capturing both local details from individual frames and long-term temporal patterns across the sequence.
-The action head and boundary head predict the action and start-end labels, respectively Nwoye et al.
-(2022).
-For a sequence length $T$ and we define the base temporal window size $64$, smallest window size $16$, largest window size $512$, average number of action segments $S$.
-To expand the receptive field, the effective window size $W_{d}$ for each layer $d$ is calculated using dilation as $W_{d}=W\cdot(1+d)$ where $W$ is the base window size, $d$ is the dilation factor ($d=0,1,2,\dots$).
-For Dual Window Sliding Attention, the effective receptive field starts at $16$ and grows up to $512$, that means $W_{d}\in[16,512]$.
-Local attention is computed within each window and then aggregated across windows.
-The temporal hierarchical attention operates at multiple scales.
-The scale $T_{s}$ is reduced iteratively $T_{s}=\frac{T}{2^{s}}$ for $s=0,1,2,\dots$, until $T_{s}\geq 512$.
-The average number of scales is approximately $\log_{2}\left(\frac{T}{S}\right)$, based on the sequence length $T$ and the average number of action segments $S$.
-At each scale $T_{s}$, global attention weights are computed using a sparse mechanism with dilation.
-For a token $x_{i}$ attending to token $x_{j}$ at scale $T_{s}$, the attention weight $\alpha_{ij}^{s}$ is
-$\alpha_{ij}^{s}=\frac{\exp\left(e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}^{s}}\exp\left(e_{ik}^{s}\right)}$ (1)
-where $e_{ij}^{s}$ is the attention score at scale $T_{s}$, $\mathcal{N}_{i}^{s}$ is the neighborhood of token $x_{i}$ at scale $T_{s}$, defined by the dilated window size $W_{d}$.
-The final attention is calculated by aggregating scores across all scales and normalizing them with a softmax function:
-$\alpha_{ij}=\frac{\exp\left(\sum_{s}e_{ij}^{s}\right)}{\sum_{k\in\mathcal{N}_{i}}\exp\left(\sum_{s}e_{ik}^{s}\right)}$ (2)
-where $\mathcal{N}_{i}=\bigcup_{s\in S}\mathcal{N}_{i}^{s}$ is the union of neighborhoods across all scales, $\alpha_{ij}$ is the final attention weight for token $x_{i}$ attending to $x_{j}$.
-At each scale $T_{s}$, the attention score $e_{ij}^{s}$ between tokens $x_{i}$ and $x_{j}$ within the neighborhood $\mathcal{N}_{i}^{s}$ is:
-$e_{ij}^{s}=\text{sim}(x_{i},x_{j}^{s})$ (3)
-where $x_{j}^{s}$ is the representation of token $x_{j}$ at scale $s$, $\text{sim}(\cdot,\cdot)$ is a similarity function (e.g., dot product).
-By restricting computations to sparse attention neighborhoods $\mathcal{N}_{i}$, the computational overhead is significantly reduced while retaining efficiency.
-The number of scales is determined byNumber of $\text{Number of scales}=\log_{2}\left(\frac{T}{S}\right).$.
-The scale reduction follows $T_{s}=\frac{T}{2^{s}},\quad s=0,1,2,\dots,\log_{2}\left(\frac{T}{S}\right),$ reducing the sequence length $T$ to approximately $\frac{T}{S}$.
-For example, if $T=1000$ and $S=4$, the segment sizes would be $(1000,500,250,125,\dots)$.
+For each query token (frame representation), attention is computed over two distinct local windows.
+Across the stack of encoder layers, one set of windows progressively expands (e.g., from a one-sided window of 16 up to 256, for an effective receptive field of 32 to 512), while the other set of windows symmetrically shrinks.
+This dynamic dual-view allows the model to capture dependencies at varying scales throughout the network depth.
+Within each sliding window, attention is dilated (e.g., with rates 0, 1, 2, …) in deeper layers.
+The dilation rates start from 0, meaning some windows, especially in earlier layers or specific attention heads, operate without dilation, focusing squarely on immediate local context.
+This allows the window to cover a larger receptive field without increasing the number of attended tokens.
+This dynamically sized windows, ensures that smaller, more focused windows are always present to capture fine-grained local details, even as other windows expand or become more dilated to see broader context.
+To complement the local/sparse sliding windows and capture broader context, we incorporate a HTA at multiple temporal scales.
+The sequence length $T_{reduced}$ is conceptually downsampled e.g., $T_{s}=T_{\text{reduced}}/2^{s}$, for scales $(s=0,1,2,\ldots)$.
+At each scale $(T_{s})$, sparse attention is computed.
+The number of scales can be determined by $\log_{2}(T_{\text{reduced}}/S_{\text{avg}})$, where $(S_{\text{avg}})$ might relate to an average segment length or a minimum sensible sequence length for global attention (e.g., until $(T_{s}\geq 512)$ as mentioned, though this was for original T).
+The attention scores from these different scales are then aggregated to form a "global" contextual representation.
+Let $(e_{ij}^{s})$ be the attention score between token $(x_{i})$ and $(x_{j})$ within its neighborhood $(\mathcal{N}_{i^{s}})$ at scale (s).
+The aggregated attention score considers information across scales:
+$\alpha{ij}=\frac{\exp\left(\sum_{s}w_{s}\cdot e_{ij}^{s}\right)}{\sum_{k\in\bigcup_{s}\mathcal{N}i^{s}}\exp\left(\sum_{s}w_{s}\cdot e{ik}^{s}\right)}$ (1)
+where$(w_{s})$ could be learnable weights or fixed, and the neighborhood $(\mathcal{N}_{i})$ spans relevant tokens across scales.
+We believe these design elements, ensure that the HSWA captures sufficient local details while effectively balancing the learning of both local and global information.
+The ablation studies presented in Section 4.3.1, especially regarding the performance of ACHDSWO (Acausal Hierarchical Dual Sliding Window with Overlap), further validate the effectiveness of this combined approach.
 3.3 Boundary-aware Segmentation Loss
@@
 Surgical videos are typically long, and the duration of the segment fluctuates significantly, as shown in Table 1.
-To address this, we propose a Boundary-Aware Segmentation Loss that integrates multiple loss functions, each targeting specific aspects of the segmentation task.
-To handle class imbalances and ensure precise classification at each timestep, we employ Focal Loss.
-This loss function down-weights well-classified examples, allowing the model to focus on harder, misclassified examples.
-This approach is particularly beneficial in datasets where certain action classes are underrepresented.
-To enhance the overlap between predicted and ground truth segments, we utilize Dice Loss derived from the Dice Coefficient:
+To address this, we propose a Boundary-Aware Segmentation Loss ($\mathcal{L}_{\text{temporal}}$ ), a unified objective function that integrates multiple specialised loss components.
+To ensure precise action classification at each timestep and handle class imbalances prevalent in surgical datasets, we employ Focal Loss $\mathcal{L}_{\text{class}}$.
+To enhance the overlap between predicted and ground truth segments, we utilise Dice Loss derived from the Dice Coefficient:
 Dice $\text{Dice Coefficient}=\frac{2\times|A\cap B|}{|A|+|B|}$ (2)
@@
 $\mathcal{L}_{\text{dice}}=1-\text{Dice Coefficient}$ (3)
-This formulation emphasizes the overlap between predictions and ground truth, making it sensitive to both false positives and false negatives.
-However, focal and dice loss primarily focuses on frame-wise classification accuracy without explicitly considering the temporal relationships between frames.
-As can be seen in figure 5, the uppermost figure shows the similarity between neighbouring frames, which are ambiguous; we smoothen the distribution by measuring centre-boundary Gaussian weighted cosine similarity depicted in the middle of the figure.
-To capture temporal dependencies and ensure smooth transitions within action segments, we introduce a loss term based on Gaussian-weighted cosine similarity.
-This loss penalizes deviations in feature representations between consecutive frames, weighted by a Gaussian function centred on action segments:
-$\mathcal{L}_{\text{sim}}=\sum_{t=1}^{T-1}G(t,t+1)\left(1-\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})\right)$ (6)
-where $G(t,t+1)$ is the Gaussian weight emphasizing temporal proximity, and $\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})$ denotes the cosine similarity between feature vectors at frames $t$ and $t+1$.
+This formulation emphasises the overlap between predictions and ground truth, making it sensitive to both false positives and false negatives.
+While Focal and Dice losses address classification and overlap, they do not explicitly model the temporal smoothness of feature representations within an action segment.
+As illustrated in Figure 5, raw frame-to-frame feature similarities can be ambiguous.
+To encourage temporal consistency and smooth transitions, $\mathcal{L}_{\text{sim}}$ penalises deviations in feature representations $\mathbf{f}_{t}$ between consecutive frames.
+This penalty is weighted by a Gaussian function $G(t,t+1)$ which is designed to peak at the temporal centre of each ground truth action segment, diminishing towards its start and end points.
+This ensures that the strongest enforcement of feature similarity occurs within the core of an action, while allowing for more natural feature evolution near the transitions to adjacent actions, promoting coherent feature sequences.
+$\mathcal{L}_{\text{sim}}=\sum_{t=1}^{T-1}G(t,t+1)\left(1-\cos(\mathbf{f}_{t},\mathbf{f}_{t+1})\right);G(t)=\exp\left(-\frac{(t-c)^{2}}{2\sigma^{2}}\right)$ (4)
+Where $c$ is the temporal centre of the segment, $\mathbf{f}_{t}$ is the feature vector of frame $t$, and $\sigma$ is a hyperparameter controlling the width of the Gaussian weighting.
+This formulation ensures that the penalty for feature dissimilarity between consecutive frames ($\mathbf{f}_{t},\mathbf{f}_{t+1}$) is highest when those frames (referenced by index $t$) are near the centre ($c$) of their action segment and lowest when they are near the boundaries.
+This aligns perfectly with the goal of encouraging strong feature similarity at the "core of an action" while allowing for more natural feature evolution near the transitions.
 To further refine boundary precision, we incorporate a Gaussian Similarity Truncated Mean Squared Error (MSE) loss.
-This loss focuses on boundary regions, applying a Gaussian weight to the MSE, which emphasizes errors near segment boundaries:
+This loss focuses on boundary regions, applying a Gaussian weight to the MSE, which emphasises errors near segment boundaries:
 $\mathcal{L}_{\text{boundary}}=\sum_{t=1}^{T}G_{\text{boundary}}(t)\cdot\min\left((\hat{b}_{t}-b_{t})^{2},\tau\right)$ (5)
-Where $G_{\text{boundary}}(t)$ is a Gaussian weight centred on the boundary, $\hat{b}_{t}$ and $b_{t}$ are the predicted and ground truth boundary positions at frame $t$, and $\tau$ is a truncation threshold to limit the influence of large errors.
+Where $G_{\text{boundary}}(t)$ is a Gaussian weight which is designed to have its highest values precisely at the annotated start and end frames of action segments (i.e., at the actual boundary points between differing actions), effectively creating high-importance zones for the loss calculation at these critical transitions.
+$\hat{b}_{t}$ and $b_{t}$ are the predicted and ground truth boundary positions at frame $t$, and $\tau$ is a truncation threshold to limit the influence of large errors.
 By integrating these loss components, the total temporal loss function is defined as:
 $\mathcal{L}_{\text{temporal}}=\alpha\cdot\mathcal{L}_{\text{class}}+\beta\cdot\mathcal{L}_{\text{dice}}+\gamma\cdot\mathcal{L}_{\text{sim}}+\delta\cdot\mathcal{L}_{\text{boundary}}$ (6)
-Where $\alpha$, $\beta$, $\gamma$, and $\delta$ are weighting factors that balance the contributions of each loss term.
+Where $\alpha$, $\beta$, $\gamma$, and $\delta$ = (1.0,0.2,0.5,0.5) were empirically determined on our validation set.
+This weighting factors balance the contributions of each loss term.
 This comprehensive loss function enables the model to effectively capture class imbalances, segment overlaps, temporal dependencies, and boundary precision, leading to improved performance in temporal action segmentation tasks.
@@
 P7 Gallbladder retraction 83 $\pm$ 56
-Table 3: Comparison of Models Across Datasets (SAR-RARP, Jigsaw, Cholec80)
-Dataset Settings Acc Prec Rec F1@10 F1@25 F1@50 Edit SAR-RARP CSAtt 0.61 0.68 0.59 0.56 0.52 0.28 0.73 CHAtt 0.73 0.77 0.71 0.58 0.54 0.40 0.75 CSFWNO 0.82 0.89 0.80 0.77 0.63 0.49 0.74 CSSWSO 0.84 0.87 0.82 0.79 0.65 0.51 0.76 CDWSO 0.87 0.85 0.81 0.85 0.68 0.55 0.79 CHFW 0.84 0.87 0.81 0.80 0.66 0.52 0.76 CHDSWO 0.88 0.89 0.80 0.84 0.72 0.58 0.80 ACSAtt 0.65 0.67 0.61 0.57 0.57 0.43 0.74 ACHAtt 0.68 0.65 0.60 0.64 0.60 0.47 0.80 ACSFWNO 0.87 0.89 0.80 0.83 0.69 0.56 0.79 ACDWSO 0.88 0.91 0.84 0.84 0.71 0.58 0.81 ACHFW 0.88 0.90 0.87 0.82 0.71 0.59 0.81 ACHSSWO 0.89 0.92 0.88 0.85 0.72 0.60 0.82 ACHDSWO 0.90 0.86 0.87 0.87 0.75 0.64 0.84 Cholec80 CSAtt 0.631 0.712 0.621 0.601 0.585 0.472 0.732 CHAtt 0.748 0.799 0.728 0.718 0.604 0.503 0.751 CSFWNO 0.84 0.821 0.83 0.81 0.695 0.583 0.762 CSSWSO 0.856 0.837 0.846 0.826 0.713 0.604 0.78 CDWSO 0.89 0.87 0.86 0.86 0.77 0.64 0.84 CHFW 0.864 0.845 0.854 0.834 0.722 0.614 0.789 CHDSWO 0.878 0.859 0.868 0.848 0.738 0.631 0.805 ACSAtt 0.686 0.767 0.676 0.656 0.547 0.542 0.741 ACHAtt 0.694 0.754 0.684 0.764 0.656 0.553 0.773 ACSFWNO 0.89 0.871 0.88 0.86 0.752 0.648 0.819 ACDWSO 0.902 0.883 0.892 0.872 0.765 0.664 0.832 ACSSWSO 0.91 0.891 0.9 0.88 0.774 0.675 0.841 ACHFW 0.906 0.887 0.896 0.876 0.77 0.67 0.837 ACHDSWO 0.918 0.899 0.908 0.888 0.783 0.686 0.85 Jigsaw CSAtt 0.742 0.823 0.732 0.712 0.598 0.488 0.705 CHAtt 0.759 0.84 0.79 0.729 0.617 0.509 0.748 CSFWNO 0.851 0.832 0.841 0.821 0.708 0.599 0.775 CSSWSO 0.867 0.848 0.857 0.837 0.726 0.620 0.793 CDWSO 0.89 0.88 0.887 0.86 0.76 0.638 0.809 CHFW 0.875 0.856 0.865 0.845 0.735 0.630 0.802 CHDSWO 0.889 0.87 0.879 0.859 0.751 0.647 0.818 ACSAtt 0.797 0.867 0.687 0.867 0.76 0.658 0.827 ACHAtt 0.905 0.886 0.895 0.875 0.769 0.669 0.836 ACSFWNO 0.901 0.882 0.891 0.871 0.765 0.664 0.832 ACDWSO 0.913 0.894 0.903 0.883 0.778 0.680 0.845 ACHFW 0.917 0.898 0.907 0.887 0.783 0.686 0.85 ACHFSWO 0.939 0.94 0.929 0.919 0.796 0.723 0.893 ACHDSWO 0.921 0.902 0.911 0.891 0.787 0.691 0.854
-Table 4: Comparison with state-of-the-art models Across Datasets (RARP, Cholec80, JIGSAW)
-Dataset Model Acc Edit F1@10 F1@25 F1@50 RARP MS-TCN Li et al.
+Table 2: Comparison with state-of-the-art models Across Datasets (SAR-RARP, Cholec80, JIGSAW)
+Dataset Model Acc Edit F1@10 F1@25 F1@50 SAR-RARP MS-TCNLi et al.
 (2020) 69.0 - 70.7 ActionCLIPPsychogyios et al.
@@
 (2021) $\boldsymbol{92.8\pm 3.3}$ $\boldsymbol{89.8\pm 6.7}$ $83.5\pm 4.2$ $70.1\pm 3.5$ $65.4\pm 1.8$ MSBATN (Ours) $86.1\pm 6.3$ $89.5\pm 6.5$ $\boldsymbol{84.4\pm 5.5}$ $\boldsymbol{76.1\pm 3.1}$ $\boldsymbol{70.7\pm 2.8}$
-Table 5: Performance comparison between Start-End Gaussian cosine and BCE Loss
-Settings Start-End Gaussian Cosine Loss Binary Cross Entropy Loss
-F1@10 F1@25 F1@50 Edit F1@10 F1@25 F1@50 Edit (Bin)
-ACDWSO 0.85 0.71 0.53 0.81 0.76 0.6 0.58 0.78
-ACHFW 0.83 0.7 0.55 0.78 0.7 0.51 0.45 0.72
-ACSSWSO 0.86 0.72 0.6 0.82 0.73 0.62 0.55 0.79
-ACHDSWO 0.87 0.74 0.63 0.84 0.77 0.64 0.53 0.8
-We extracted every frame (256x 256x3) from the video sequences and used the Dinov2-Reg feature extractor to transform these frames into a 2048-dimensional feature space.
-These features serve as inputs to our encoder-decoder architecture.
+Figure 6: Qualitative action segmentation comparison.
+(a) Ground Truth (GT), where each colour label represents an action.
+The rectangular boxes explicitly highlight actions for detailed comparison of actions’ start/end.
+(b) ASRF Ishikawa et al.
+(2021) and (c) Asformer Yi et al.
+(2021) show deviations in action start/end points (highlighted in red rectangular box).
+(d) The proposed MSBATN model demonstrates better alignment with GT (highlighted in green).
+Segments are zoomed from the full video timeline below.
 The encoder ingests inputs of shape (T, 2048) and processes them through a dilated acausal TCN, effectively reducing the dimensionality to (T,256) while capturing temporal dependencies over varying scales.
-These embeddings are further refined through ten layers of multi-head attention mechanisms (with 8 heads), incorporating advanced techniques such as sliding window attention for local dependencies and hierarchical attention for multi-scale temporal modelling.
+These embeddings are further refined through ten layers of multi-head sparse attention mechanisms (with 8 heads), incorporating advanced techniques such as sliding window attention for local dependencies and hierarchical attention for multi-scale temporal modelling.
 This combination enables the encoder to learn both fine-grained and long-range temporal patterns.
@@
 Each decoder shares a similar structure to the encoder but focuses on progressively refining predictions with higher accuracy and temporal smoothness.
-For optimization, we utilized the Adam optimizer along with the softmax activation function to ensure stable training.
+For optimisation, we utilised the Adam optimiser along with the softmax activation function to ensure stable training.
 In contrast to some existing approaches Yi et al.
@@
 Our boundary-aware loss function is designed to enhance the model’s segmentation and transition detection capabilities.
-It combines a weighted Dice loss, which penalizes the model based on Intersection over Union (IoU), with a Gaussian Cosine Similarity loss that emphasizes smooth transitions by modelling centre-to-boundary similarity.
+It combines a weighted Dice loss, which penalises the model based on Intersection over Union (IoU), with a Gaussian Cosine Similarity loss that emphasises smooth transitions by modelling centre-to-boundary similarity.
 This composite loss function ensures robust learning of temporal dynamics and accurate identification of action boundaries.
@@
 A higher Edit Score indicates a closer match between the predicted and actual sequences, reflecting the model’s proficiency in capturing the correct order and transitions of actions.
-This metric is less sensitive to minor temporal shifts but penalizes over-segmentation and out-of-order predictions.
+This metric is less sensitive to minor temporal shifts but penalises over-segmentation and out-of-order predictions.
 Calculate the Levenshtein distance $D$ between the predicted and ground truth sequences.
 The Edit Score is then computed as Edit $\text{Edit Score}=1-\frac{D}{\max(L_{p},L_{g})}$ where $L_{p}$ and $L_{g}$ are the lengths of the predicted and ground truth sequences, respectively.
-The Segmental F1 Score is sensitive to over-segmentation and under-segmentation errors, penalizing models that divide actions into too many segments or fewer segments.
-In contrast, the Edit Score penalizes incorrect action order and over-segmentation but is more tolerant of slight temporal misalignments.
+The Segmental F1 Score is sensitive to over-segmentation and under-segmentation errors, penalising models that divide actions into too many segments or fewer segments.
+In contrast, the Edit Score penalises incorrect action order and over-segmentation but is more tolerant of slight temporal misalignments.
 The Segmental F1 Score evaluates the spatial and temporal precision of action boundaries, whereas the Edit Score assesses the structural and sequential accuracy of the predicted action sequence.
-4.2 Results Analysis
-Our Multi-staged Baundary aware network has three main differences form Asformer, First we employed acausal convolution layer in the initial layers of encoders and decoders.
+4.2 Results
+Our Multi-staged Boundary-aware network has three main differences from Asformer.
+First, we employed acausal convolution layer in the initial layers of encoders and decoders.
 Secondly, we used hierarchical sparse multi-head window attention.
 Finally, we calculated centre sampled Boundary Aware Loss leverages the strengths of both cosine similarity and Gaussian functions, making it effective for segmentation tasks.
-We conducted the following experiments to demonstrate the effects of Causal and Acausal convolution along with hierarchical dual or single sliding window attentions.
-4.2.1 Effect of Convolution and sliding window attention
-We assessed the effectiveness of dilated causal and acausal convolutions, along with various attention settings during training.
-As illustrated in Table 3, the causal settings include: - Causal Self-Attention (CSAtt) - Causal Hierarchical Attention (CHAtt) - Causal Sliding Fixed Window with No Overlap (CSFWNO) - Causal Sliding Single Window with Overlap (CSSWO) - Causal Dual Sliding Window with Overlap (CDWSO) - Causal Hierarchical Fixed Window with Overlap (CHFWO) - Causal Hierarchical Dual Sliding Window with Overlap (CHDSWO)
-In addition, we incorporated acausal convolutions, which include: - Acausal Self-Attention (ACSAtt) - Acausal Hierarchical Attention (ACHAtt) - Acausal Sliding Fixed Window with No Overlap (ACSFWNO) - Acausal Sliding Single Window with Overlap (ACSSWO) - Acausal Dual Sliding Window with Overlap (ACDWSO) - Acausal Hierarchical Fixed Window with Overlap (ACHFWO) - Acausal Hierarchical Dual Sliding Window with Overlap (ACHDSWO) Through ongoing evaluation of these settings, we concluded that for longer sequences, Dual Sliding Overlapping Window Attention outperformed Single Sliding Overlapping Window.
-Conversely, for shorter action sequences in the Jigsaw dataset, the Single Sliding Window demonstrated superior performance.
+As depicted in Table 2 our model outperforms other models in F1 scores and achieves comparable results in Acc and Edit with state of the art.
+While Accuracy and Edit Scores provide valuable insights into frame-level precision and strict sequential ordering, they are highly sensitive to slight temporal misalignment at segment boundaries.
+In contrast, the F1@k metric offers greater flexibility for boundary localisation (±$(\pm k\%)$ making it more robust to accurately segment long actions, even with minor shifts in timing.
+Figure 6 illustrates this qualitatively, showcasing MSBATN’s superior performance in accurately defining crucial action boundaries compared to other methods like ASRF and Asformer.
+Table 5 shows MSBATN is computationally efficient in comparison with other boundary aware model like ASRF Ishikawa et al.
+(2021).
+Table 3: Comparison of different settings Across Datasets (SAR-RARP, Jigsaw, Cholec80)
+Dataset Settings Acc Prec Rec F1@10 F1@25 F1@50 Edit SAR-RARP CSAtt 0.61 0.68 0.59 0.56 0.52 0.28 0.73 CHAtt 0.73 0.77 0.71 0.58 0.54 0.40 0.75 CSFWNO 0.82 0.89 0.80 0.77 0.63 0.49 0.74 CSSWSO 0.84 0.87 0.82 0.79 0.65 0.51 0.76 CDWSO 0.87 0.85 0.81 0.85 0.68 0.55 0.79 CHFW 0.84 0.87 0.81 0.80 0.66 0.52 0.76 CHDSWO 0.88 0.89 0.80 0.84 0.72 0.58 0.80 ACSAtt 0.65 0.67 0.61 0.57 0.57 0.43 0.74 ACHAtt 0.68 0.65 0.60 0.64 0.60 0.47 0.80 ACSFWNO 0.87 0.89 0.80 0.83 0.69 0.56 0.79 ACDWSO 0.88 0.91 0.84 0.84 0.71 0.58 0.81 ACHFW 0.88 0.90 0.87 0.82 0.71 0.59 0.81 ACHSSWO 0.89 0.92 0.88 0.85 0.72 0.60 0.82 ACHDSWO 0.90 0.86 0.87 0.87 0.75 0.64 0.84 Cholec80 CSAtt 0.631 0.712 0.621 0.601 0.585 0.472 0.732 CHAtt 0.748 0.799 0.728 0.718 0.604 0.503 0.751 CSFWNO 0.84 0.821 0.83 0.81 0.695 0.583 0.762 CSSWSO 0.856 0.837 0.846 0.826 0.713 0.604 0.78 CDWSO 0.89 0.87 0.86 0.86 0.77 0.64 0.84 CHFW 0.864 0.845 0.854 0.834 0.722 0.614 0.789 CHDSWO 0.878 0.859 0.868 0.848 0.738 0.631 0.805 ACSAtt 0.686 0.767 0.676 0.656 0.547 0.542 0.741 ACHAtt 0.694 0.754 0.684 0.764 0.656 0.553 0.773 ACSFWNO 0.89 0.871 0.88 0.86 0.752 0.648 0.819 ACDWSO 0.902 0.883 0.892 0.872 0.765 0.664 0.832 ACSSWSO 0.91 0.891 0.9 0.88 0.774 0.675 0.841 ACHFW 0.906 0.887 0.896 0.876 0.77 0.67 0.837 ACHDSWO 0.918 0.899 0.908 0.888 0.783 0.686 0.85 JIGSAW CSAtt 0.742 0.823 0.732 0.712 0.598 0.488 0.705 CHAtt 0.759 0.84 0.79 0.729 0.617 0.509 0.748 CSFWNO 0.851 0.832 0.841 0.821 0.708 0.599 0.775 CSSWSO 0.867 0.848 0.857 0.837 0.726 0.620 0.793 CDWSO 0.89 0.88 0.887 0.86 0.76 0.638 0.809 CHFW 0.875 0.856 0.865 0.845 0.735 0.630 0.802 CHDSWO 0.889 0.87 0.879 0.859 0.751 0.647 0.818 ACSAtt 0.797 0.867 0.687 0.867 0.76 0.658 0.827 ACHAtt 0.905 0.886 0.895 0.875 0.769 0.669 0.836 ACSFWNO 0.901 0.882 0.891 0.871 0.765 0.664 0.832 ACDWSO 0.913 0.894 0.903 0.883 0.778 0.680 0.845 ACHFW 0.917 0.898 0.907 0.887 0.783 0.686 0.85 ACHFSWO 0.939 0.94 0.929 0.919 0.796 0.723 0.893 ACHDSWO 0.921 0.902 0.911 0.891 0.787 0.691 0.854
+4.3 Ablation studies
+We conducted an ablation study to demonstrate the effectiveness of the components in our architecture.
+Our configuration, which includes Acausal Convolution with Hierarchical Dual Sliding with Overlap, as shown in Table 3, outperforms other settings in long action segmentation.
+Additionally, we evaluated the computational complexity, as detailed in Table 5.
+Furthermore, we compared the effectiveness of the loss functions, as presented in Table 4.
+4.3.1 Effect of Convolution and sliding window attention and overlapping
+We assessed the effectiveness of convolutions, along with various attention settings during training.
+As illustrated in Table 3, the causal settings include: - Causal Self-Attention (CSAtt) - Causal Hierarchical Attention (CHAtt) - Causal Fixed Sliding Window with No Overlap (CFSWNO) - Causal Single Sliding Window with Overlap (CSSWO) - Causal Dual Sliding Window with Overlap (CDWSO) - Causal Hierarchical Fixed Window with Overlap (CHFWO) - Causal Hierarchical Dual Sliding Window with Overlap (CHDSWO) In addition, we incorporated acausal convolutions, which include: - Acausal Self-Attention (ACSAtt) - Acausal Hierarchical Attention (ACHAtt) - Acausal Fixed Sliding Window with No Overlap (ACFSWNO) - Acausal Single Sliding Window with Overlap (ACSSWO) - Acausal Dual Sliding Window with Overlap (ACDWSO) - Acausal Hierarchical Fixed Window with Overlap (ACHFWO) - Acausal Hierarchical Dual Sliding Window with Overlap (ACHDSWO)
 Table 3 highlights the effectiveness of proposed hierarchical sliding window attention with causal and acausal convolution.
-Settings with acausal convolution performed better than causal convolution as causal convolution restricts information flow to the present and past, ensuring that no future data influences the current output.
+Settings with acausal convolution performed better than causal convolution, as causal convolution restricts information flow to the present and past, ensuring that no future data influences the current output.
 In contrast, acausal convolution permits the output to depend on past, present, and future input values, which is useful during training.
@@
 Hierarchical Sliding Window Sparse Attention focuses on local neighbourhoods dynamically, capturing essential local patterns and structures.
-Stacking layers of such windowed attention results in a large receptive field, where top layers have access to all input locations and can build representations that incorporate information across the entire input.
+Stacking layers of such windowed attention results in a large receptive field, where the top layers have access to all input locations and can build representations that incorporate information across the entire input.
 Single-Sliding Window Attention performs well on the Jigsaw dataset used in surgical training, where samples are shorter and backgrounds are simpler than in real-world datasets like SAR-RARP and Cholec80.
@@
 Meanwhile, Dual-Sliding Window Attention is better at capturing both short- and long-range dependencies, making it suitable for handling diverse action durations and transitions.
+Table 4: Performance comparison between Start-End Gaussian cosine and BCE Loss
+Settings Start-End Gaussian Cosine Loss Binary Cross Entropy Loss
+F1@10 F1@25 F1@50 Edit F1@10 F1@25 F1@50 Edit (Bin)
+ACDWSO 0.85 0.71 0.53 0.81 0.76 0.6 0.58 0.78
+ACHFW 0.83 0.7 0.55 0.78 0.7 0.51 0.45 0.72
+ACSSWSO 0.86 0.72 0.6 0.82 0.73 0.62 0.55 0.79
+ACHDSWO 0.87 0.74 0.63 0.84 0.77 0.64 0.53 0.8
+Table 5: Impact of HSWA Components (DSWA and HTA) and GFLOPs and parameters.
+Model Acc F1@10 GFLOPs Parameters
+Transformer Vaswani et al. (2017) 65.3 57.8 204.47 20.5181 M
+ASRF Ishikawa et al. (2021) 87.6 80.3 136.42 13.6686 M
+NETE Yi et al. (2022) 87.9 79.4 131.10 13.1454 M
+Asformer Yi et al. (2021) 88.5 84.5 101.03 10.0027 M
+DSWA 86.8 82.6 101.31 7.9750 M
+HTA 84.4 76.2 76.78 7.8594 M
+HSWA(DSWA+HTA) 88.7 84.5 104.53 11.9450 M
 4.3.2 Effect of Start-End Boundary Aware Loss Function
@@
 Additionally, binary boundary classification may lead to instability during training due to substantial variations in loss values both within and between classes.
-The results of the proposed boundary-aware loss function show that our proposed approach effectively handled these challenges by capitalizing on the distribution of similarity scores between frames generated by decoders.
-Table 5 demonstrates the effectiveness of our method, which incorporates actual convolution and sliding window attention settings, delivering improved results across the five metrics outlined in the previous section compared to the binary-boundary-based approach.
+The results of the proposed boundary-aware loss function show that our proposed approach effectively handled these challenges by capitalising on the distribution of similarity scores between frames generated by decoders.
+Table 4 demonstrates the effectiveness of our method, which incorporates acausal convolution and sliding window attention settings, delivering improved results across the five metrics outlined in the previous section compared to the binary-boundary-based approach.
 5 Conclusion and Feature Work
@@
 We found that selecting the right attention module can be difficult, and neighbouring frames often lead to ambiguous results, causing over- and under-segmentation in frame-wise segmentation.
-Segment-wise refinement is more effective in improving understanding of surgical activities.
+Segment-wise refinement is more effective in improving the understanding of surgical activities.
 Additionally, binary label boundaries are often inadequate because the start, centre, and end of actions vary by individual expertise, leading to poor performance.
-Our proposed framework addresses these issues with hierarchical sliding window attention for global-local context and includes a novel Boundary Segmentation Loss to better recognize central and boundary points.
+Our proposed framework addresses these issues with hierarchical sliding window attention for global-local context and includes a novel Boundary Aware Segmentation Loss to better recognise central and boundary points.
 Our boundary-aware segmentation approach outperforms traditional binary detection methods.
@@
 Furthermore, integrating instructional prompts to enhance and refine the learning process represents another promising avenue for future investigation.
+[…]
+Declaration of generative AI and AI-assisted technologies in the writing process
+During the preparation of this work, the author(s) used Grammarly and ChatGPT in order to refine the writing, check grammar, and improve overall readability.
+After using these tools, the author(s) reviewed and edited the content as needed and take(s) full responsibility for the content of the publication.